### v2

# What Do You Need for Diverse Trajectory Composition in Diffusion Planning?

Abstract
In planning, stitching is an ability of algorithms to piece together sub-trajectories of data they are trained on to generate new and diverse behaviours. While stitching has historically been a strength of offline reinforcement learning, recent generative behavioural cloning (BC) methods have also shown proficiency at stitching. However, the main factors behind this are poorly understood, hindering the development of new algorithms that can reliably stitch. Focusing on diffusion planners trained via imitation, we find two properties are key contributors for composition: shift equivariance and local receptiveness. We use these two properties to explain architecture, data, and inference choices in existing generative BC methods based on diffusion planning, including replanning frequency, data augmentation, and data scaling. Experimental comparisons show that (1) while locality is more important than shift equivariance in creating a diffusion planner capable of composition, both are crucial (2) enabling these properties through relatively simple architecture choices can be competitive with more computationally expensive methods such as replanning or scaling data, and (3) simple inpainting-based guidance can guide architecturally compositional models to enable generalization in goal-conditioned settings.
1 Introduction
The recent resurgence in generative diffusion models [58, 30] has inspired a new wave of work investigating the potential for these models to act as planners for robotic agents [34, 13]. One surprising ability these diffusion planning models have is compositionality: the ability to take component elements seen in samples of their dataset and compose them in new, kinematically feasible ways [20, 44].
One form of compositionality especially important in diffusion planning is stitching, where some diffusion planning methods are capable of composing sub-trajectories into new trajectories distinct from the training data. Historically, stitching has been seen as one of the primary benefits of temporal difference (TD) learning [38], both in offline and online RL. Planning with diffusion models thus potentially offers some of the stitching benefits of TD learning without needing to handle RL-specific challenges.
Despite their recent success in vision, when and why diffusion models compose in the planning domain is not well understood. Contrary to toy experiments in prior works [1, 34], the convolutional neural network (CNN)-based backbones commonly used in diffusion planners often struggle with composition [9], instead regressing to memorization of training examples [2].
This gap between the promise of diffusion planners to enable stitching and reality that they frequently struggle with composition is crucial to address, as diffusion offers an appealing alternative to other stitching approaches. RL suffers from many challenges when stitching such as the deadly triad [33, 22, 64] and slow temporal credit assignment [12], while non-generative BC stitching approaches require nettlesome data augmentation [26]. The capability to sample a diverse set of new trajectories at test time is important property for planners broadly. Reasons for this include being able to select solutions from a candidate set [69], encouraging multi-modal behaviour [40], or generating diverse data to train imitation learning or offline RL algorithms [38].
To address this gap, we analyze a variety of approaches that induce composition in diffusion planners. We find that there are two key ingredients: local receptiveness [52] during generation, and shift equivariance of states in a trajectory [36]. Through experiments in navigation environments, we show that although most current diffusion planning models do not possess this compositional capacity natively, it can be induced through architecture changes, inference choices, data augmentation strategies, or data scaling without needing dynamic programming or TD-learning. We further show that the diversity of trajectories this compositionality enables can be selectively sampled from using inpainting, allowing the use of diffusion planners only trained on imitation as goal-conditioned planners, similar to offline goal-conditioned RL.
Our primary contributions are as follows:
Showing that local attention and shift equivariance, previously shown to mechanistically explain creativity and generalization in image diffusion models [36], also function as ingredients for composition in diffusion planners. However, unlike the image diffusion case, they are not the only factors: the inference method used to sample trajectories from diffusion planners also plays a significant role.
Demonstrating that when induced through architecture changes, the stitching capabilities of whole-sequence diffusion can be effectively guided through simple inpainting to achieve strong goal-conditioned performance on unseen start-goal pairs.
Our work provides more clarity behind what components are necessary to guide diffusion planners to compose trajectories at test-time successfully. We hope this will inspire future algorithmic and architectural developments in compositional diffusion planners.
2 Related Work
Diffusion Models: Diffusion models [61, 30] offer an efficient way train models that can sample from complex distributions. Diffusion models train from data points drawn from some target distribution by taking a sample and iteratively adding Gaussian noise, such that , called the forward diffusion process. The diffusion model learns the reverse diffusion process, where some model (usually a neural network) parameterized by is trained to learn by minimizing the MSE loss between the model’s output and the reverse diffusion step: . Sampling from the new distribution is performed by first sampling from the Gaussian and using the model to reverse-sample the diffusion process: [50, 7].
Diffusion Planning: Introduced by [34], diffusion planners model trajectories with diffusion. A diffusion model predicts an upcoming state or state-action trajectory horizon time-steps long, conditioned on a starting state or state memory of length , in an environment with max steps. From the generated plan, actions are executed for a fixed horizon before generating a new trajectory. In this work, we are mostly interested in the whole sequence planning domain, where , and no replanning occurs. Diffusion planning should not be confused with the diffusion policy [13] formulation where the diffusion model generates a short sequence of actions with high reactivity from frequent re-generation of actions. We elaborate on this distinction further in Appendix A.6.3.
Stitching in Imitation Learning: While stitching is often viewed as a strength of TD learning [38], other work analyzes stitching in imitation learning [8, 28]. We were inspired by [26], which analyzes stitching through a compositional generalization perspective. They revealed that applying supervised-learning as a form of offline RL [21] fails to stitch, requiring training data augmentation which manually stitches together trajectories. However, their result does not directly translate to ours because the properties we analyze (local receptiveness and shift equivariance) both enforce a form of regularization that creates composition, even when such behaviour is not explicitly built into the data.
Compositionality in Diffusion Planners: Many approaches exist to create compositionality in diffusion planners. One common thread is performing hierarchical planning, including using diffusion to stitch existing dataset trajectories [41], generating subsequences jointly conditioned on each-other [48], or using a hierarchical cascade of models to separately create high-level plans and low-level trajectories [10, 42]. Our work differs from these by mainly targeting whole-sequence compositionality, where composition does not occur from explicit stitching but from implicit model behaviour, similar to image diffusion. 111See Appendix A.6.2 for more discussion.
Several other works have observed that Diffusion models often produce stitched trajectories when frequently replanning during inference [9] and limiting the history the next generation is conditioned on [62]. Our work attempts to formalize these qualitative observations with more robust experiments and explain it under a single conceptual framework (locality for short history and shift equivariance for the replanning process) that also extends to architectural or training choices.
Our focus on shift equivariance and locality was inspired by prior works which observed that these properties explain much of the compositional behaviour of image diffusion models [36, 52]. We extend their work in a few significant ways. One, we connect their image-diffusion observations to sequential decision-making by connecting these properties that enable composition to other methods commonly used to achieve stitching in IL. Two, we show that common diffusion architectures fail to exhibit these properties, preventing stitching from occurring, especially at small data sizes. Three, we analyze the effect each independently has on compositionality. Finally, we show that common guidance strategies enable strong base-model compositionality to be directed towards certain goals, which is particularly important when using generative models for offline goal-conditioned RL.
3 Methodology
3.1 Local Receptiveness and shift equivariance in Score-Based Diffusion Models
To unify prior explanations for compositionality in diffusion planners, we rely on two properties previously described for diffusion models in vision [36, 52].
The first is locality, which means the size of the region around an individual state in a trajectory that influences the next denoising step for that state is constrained to a small region. A local model only attends to a small patch around part of a sample when denoising, but through multiple denoising steps, information can propagate across a generated sample.
Definition 1: Locally Receptive Field. Assume we are interested in generating trjectories , and have a generative trajectory model whose output estimates the denoising gradient . We say that has a field locally receptive to a state neighbourhood if for all trajectories and state positions within those trajectories , .
The second is positional equivariance, which requires that the model not use the absolute position of a state in the sequence as a condition during the denoising process. Convolutional neural networks (CNNs) are typically assumed to have this property, as convolutions by design shift a shared weight kernel and apply it to local patches in a sequence. However, downsampling and pooling operators common in CNNs break this property [70, 6].
Definition 2: Shift equivariance. A trajectory score model is shift equivariant if for all shifts 222By shift equivariant we mean to discrete shifts and not continuous translations. See [49]. on trajectories which uniformly change the position of states within the trajectory, . 333In theory, circular padding is required for a CNN to have full shift equivariance[36], but in practice this is uncommon.
3.2 Composition in Diffusion Planners
Our main insight is that these notions of locality and shift equivariance have analogies to other strategies used to increase composition in diffusion planners.
Increasing Compositionality by Replanning: The first strategy we address is replanning, specifically replanning with a limited history. A trajectory planning model replans when it generates a trajectory of length , executes some action horizon , and then generates a new trajectory starting at the current state . Because a diffusion planner can be queried to generate plans with a variable length, is often set to some number large enough to ensure temporal consistency far enough into the future to do effective planning. Additionally, a history of past states to condition on with length must be decided.
Formally, given a generative trajectory model that generates trajectories of length conditioned on a history of length , each state in the trajectory will only be directly attentive to a local field of size , i.e, .
It is important to note that locality through replanning and locality in the reverse-diffusion process are not identical, as the reverse diffusion process being several steps long means influence can be spread along a long trajectory through many local denoising steps. We show this visually in Figure 3. However, the similar mechanisms highlight how increasing the influence of nearby states in a given trajectory is important for composition.
Replanning also ensures a form of shift equivariance when replanning with a fixed amount of the past trajectory kept as conditioning for the future trajectory. Diffusion planners can condition their plans on past state observations in two ways: either by including it as an explicit condition to the diffusion network, usually combined with classifier-free guidance (CFG) [31], or by inpainting states in the new generated trajectory. Full memory with inpainting does not guarantee shift equivariance, as each state in the trajectory can end up only being drawn to the same position in the trajectory that they were seen in during training. However, replanning with a smaller history shifts the viewed positions of a particular segment of the trajectory over by steps. Paradoxically, this means models with strong replanning capability must already have some form of shift equivariance.
Increasing Compositionality by Training Augmentation: The next strategy to increase compositionality we address is training augmentation, where the data is modified during training to give the model approximate learned equivariance to some transformation [3, 59]. The most common approach to encourage stitching is manually stitching together trajectories in the training set randomly by finding similar states in different trajectories[26, 8, 28, 39]. This augmentation attempts to breaks the corelation between states far away in a trajectory that prevents stitching. However, this strategy requires tuning the distance metric that determines when segments are close, and is potentially infeasible if the combinatorial set of all legal stitches is large enough to make augmenting every possible trajectory expensive.
We instead focus on augmenting the data to provide shift equivariance. Specifically, we examine positional augmentation, where the position of specific states within a trajectory are randomized during training such that the model cannot overfit to a state’s position. We implement this by sampling a random state from a training trajectory, removing the states before that state, and then padding the end of the trajectory with the final state until it is the same length as the original trajectory, such that the trajectory is randomly "shifted" by a number of states.444This assumes that the environment can be kept at the same state by the agent, which is true of our environments. Alternatives include masking out the remainder of the trajectory or allowing for mixed-length trajectories during training.
Formally, given a training trajectory with states , we sample a position from the uniform distribution . To create our new augmented trajectory , we shift the states in by positions to the left, such that for state indexes , . For the remainder of the states , we set pad the states with the last state in the original trajectory: .This modified trajectory is then used when training our planner.
Increasing Compositionality via Architecture Design Choices: The next strategy we address are architecture modifications. While equivariance can be learned through data augmentation, another popular strategy is to simply build equivariance to certain transformations through architecture changes[14, 15], especially in limited data settings.
Surprisingly, we found that the common 1D U-Net proposed in [34] has a globally receptive field and is not equivariant to shifts, which we show in Appendix A.4.1. To easily analyze the effect that locality and shift equivariance have on composition, we replace the U-Net with a network which we call Eq-Net. Eq-Net simply consists of stacked 1D convolutions without pooling or down-sampling layers, which guarantees equivariance to shifts. We also decrease the size of the kernels in Eq-Net to guarantee locality. We defer further discussion and explanation to Appendix A.4.2.
To test the effects of having a network with local receptiveness but not shift equivariance, we add positional encoding to Eq-Net with sinusoidal embeddings [66], adding each time positions’ embedding channel-wise after the first convolutional layer. See Appendix A.4.5 for details.
Increasing Compositionality via Scaling Training Data: The last strategy we address is scaling training data, which has been frequently observed to increase diffusion model generalization [68] and compositionality [53]. Importantly, this increase in compositionality does not come from memorizing more of the legal space of generations, but from learning the underlying rules governing composition in a particular area. Specifically, [23] finds that as data scales, progressively higher-level grammar rules are learned, starting with low level (local) consistency rules.
Our work is specific to sequential decision making, where data is often sparse. For instance, the toy vision dataset CIFAR-10 [37] has samples, and the more modern ImageNet [17] has . By contrast, most of the environments in the modern offline RL OGBench dataset [54] have between and episodes per environment. Similarly, most of the tasks in [13] use or less trajectories. Given that most sequential decision making datasets have orders of magnitude fewer samples, considering other approaches to generalization becomes more important than relying on data scaling.
Explaining Composition in Other Diffusion Planners: Our goal in examining the factors that lead to composition is to help explain why prior works compose (or fail to) so that future diffusion planning practitioners know which choices are important if their domain requires compositionality. Due to space constraints we include a thorough discussion of other works in Appendix A.6.3. Our main finding is that the majority of existing works achieve locality and shift equivariance either through frequent replanning with short-history conditioning [62, 9, 13] or some form of hierarchical decomposition where trajectories are generated explicitly as sub-trajectories and then composed manually [10, 48, 42]. Surprisingly, achieving compositionality through architecture choices alone is rare, with our work to our knowledge being the first to take this approach in the planning domain.
4 Experiments
Our hypothesis is that local receptiveness and positional equivariance are key ingredients in explaining compositionality in diffusion planning, and that different ways of incorporating each (training augmentations, replanning, architecture choices, etc.) should increase compositionality. However, it is not known which of the two properties is more important, whether both are needed, or which way of inducting the two properties increases compositionality the most effectively.
To test our hypothesis that local receptiveness and shift equivariance are sufficient for composition in diffusion planners and answer our open questions, we train a variety of models on toy environments designed to test generated sequence diversity and multi-goal completion.555Our code can be found here: https://github.com/rvl-lab-utoronto/diffusion-stitching
Environment Description: Our main environment is a navigation task where an agent must navigate from one point in a grid to another while avoiding obstacles. Figure 1 shows examples of our grid environment, including points navigating to and from different locations in the grid. Additionally, to test training data scaling, we use a separate navigation environment where an agent moves between ten points in euclidean space in different sequences of points. More details are deferred to Appendix A.3.
We chose toy navigation environments over complex benchmarks for a few of reasons. First, popular datasets which claim to test stitching either have testing start-goal pairs contained in the training set [24]666Shown by [26]. or a very limited number of start and goal states [54]. Additionally, existing datasets do not report what subsequences are present in the training data, making it difficult to predict stitching behaviour, whereas we can control this in our environment. Finally, the grid layout of our environment makes quantifying diversity easy, as each stitched trajectory will be topologically distinct [57], allowing us to perform careful analysis of a given model’s compositional ability.
Data Collection for Diffusion Model Training: We look at the diversity of trajectories generated by diffusion planners in two separate planning settings, and collect two seperate datasets in our grid environment.
The first is the fixed-endpoint setting, where the goal of the task is to reach a fixed end point from a fixed start point. Every training trajectory is optimal for the task, but the demonstration data includes multi-modal behaviour, as individual trajectories reach the end point via a different route. All trajectories have the same starting and ending state but take different paths through the grid, and the diffusion planner is tasked with generating new paths from the start to the goal composed of sub-trajectories seen in training in new configurations.
The second is the end-point conditional setting, where a wider range of trajectories provide sub-trajectories that need to be composed to meet novel start-goal pairs. In our environment, the goal is defined as reaching a goal state from a start state, where each intersection point along the edge of the grid is a legitimate start or goal state. To be precise, we attempt to learn a model that can sample from the distribution . We collect data that includes diverse grid navigation to give the model ample sub-trajectories to compose, while being careful to keep some start-goal pairs unseen. We defer details to Appendix A.3.
Algorithm Setup: For our experiments we use a 1D U-Net Diffusion Planner as the base model. We use CleanDiffuser’s [18] implementation and hyperparameters. One change we make is only predicting future states and using an inverse dynamics model to recover actions [1]. 777See Table 2 for details on hyperparameters and inverse dynamics..
5 Results
Here we report our findings on the unconditional and conditional environments, with the environments, data, and algorithms described in Section 4.
In our results, we count diversity by examining topologically distinct trajectories. Practically, each trajectory can split the obstacles into two groups: those on its left side, and on its right; two trajectories with different obstacle splits are topologically unique.
5.1 Unconditional Environment Results
To quantify each model’s ability to compose in the unconditional environment, we took our final trained models and sampled them at a range of sample counts. From these samples, we counted what percentage of the possible (but unseen in training) trajectory space each model’s generations covered. By giving us a sense of how a model’s coverage scales with the number of generations, this gives a sense of how capable each model’s composition abilities are without solely relying on a model’s behaviour at very high sampling counts.888In our unconditional results, we filter to only analyze successful trajectories (ones that successfully reached the end state without hitting an obstacle) but all methods achieved very high success rates. Full completion rates are reported in Appendix A.2.1.
Figure 4 shows our results for the unconditional navigation environment. Figure 4(a) shows a comparison of the different whole-sequence composition strategies. The base U-Net fails to ever compose on our small dataset, only reproducing examples seen in training. Our model with approximate positional equivariance (granted through training augmentation) performs better, but fails to extrapolate, not covering a significant proportion of the unseen trajectory space even at very high sampling rates. Our equivariant network with added positional encoding (making the model local, but not positionally equivariant) performs much better, indicating locality is a more important factor in encouraging composition than positional equivariance. Lastly, our model that is both locally receptive and shift equivariant (Eq-Net) performs the best, exhibiting very strong whole-sequence composition from architecture choices alone, without utilizing data scaling, augmentation, or replanning.
Figure 4(b) analyzes replanning, where we progressively increase the number of times the model replans during a trajectory, and analyze the resulting diversity of generations over independent generations. For our replanning model we used our U-Net trained with positional data augmentation, as we found it necessary to enable the model to generate trajectories from diverse start positions. Assume we wish to replan times in a task long. When replanning, we keep only the current state ( using our prior notation), generate a new trajectory starting at that point, execute steps, and then repeat until the goal is reached. Figure 4(b) shows that as expected, as , the diversity of generations increases. To compare this to our architectural strategy, we compare it against the best model from Figure 4. To directly compare the computational cost associated with replanning, we compare each agent that replans times against Eq-Net that gets times as many parallel samples to generate from, such that each point queries the diffusion model an equal amount of times. For instance, at , the replanning model over independent samples is queried times, as each trajectory requires queries, so to have a fair comparison we query Eq-Net for independent samples.
Predictably, our results show that more frequent replanning indeed leads to more trajectory diversity. Interestingly, they also show that the extra budget of diffusion model queries required during replanning might be better spent generating more samples from a high-diversity whole-trajectory model, as Eq-Net performs similarly or slightly better on all samples.
Lastly, to validate that our results are not influenced by regularization induced by under-training of our diffusion planner, we analyze how each model’s compositionality evolves throughout training. We find that while early stopping helps compositionality (perhaps acting as a form of regularization that prevents overfitting to the training samples as has been observed in other work [5]), models with stronger inductive biases preserve their compositional abilities as they trained for longer. See Appendix A.5.1 for details.
5.2 Conditional Environment Results
In the conditional environment described in Section 4, the goal of the diffusion planner is to take unstructured trajectories in our environment and be able to consistently produce trajectories connecting an arbitrary start and goal location. To ensure that stitching is required at inference time, we examine every pair of edge intersections and filter out any pair where the starting pair and goal pair appear together sequentially, resulting in test-time pairs. We run our whole-sequence planners on each pair times, calculate an average completion rate and number of unique successful plans (where uniqueness is defined as topological distinctness), and take the average over sample runs.
To guide the model to a specific goal, we use inpainting like in prior diffusion works [47, 34, 61]. We found that repeating the inpainted goal state significantly improved goal-reaching performance, so we repeat the goal state times at the end of the trajectory before diffusing the intermediate states.
Our results are shown in Figure 5. The model that is both local and positionally equivariant far outperforms all other models, achieving an average completion rate, roughly higher than the model with only locality but not positional equivariance. Only having positional equivariance did not substantially help the model in this scenario, with the U-Net with data augmentation only achieving a non-zero result on one of the testing pairs and only a small amount of the time. Lastly, like in the unconditional case, the default U-Net fails to perform stitching entirely if it is not locally receptive or shift equivariant.
To visualize the difference in composition for each model, we plotted the unique successful runs from random generations from each algorithm on one of the hold-out start-goal pairs. Figure 6 shows the relative diversity of each model: the base U-Net fails to ever generate a successful trajectory, the U-Net made approximately positionally equivariant with data augmentation can produce a single candidate successful trajectory, the Eq-Net with positional equivariance removed produces a handful of plans, and the full Eq-Net produces a large diversity of successful trajectories.
5.3 Data Scaling Results
Lastly, we investigate the role of data in inducing composition. We use a different environment in these experiments, consisting of a navigation task where a point must sequentially navigate between five different points of a decagon. A compositional planner should be able to stitch sub-trajectories from individual points to other points. We train models with data ranging from from to data points, where points is close to the number of possible combinations.
Figure 7 shows that even models without strong inductive biases can learn to compose at large enough data sizes. Past a certain data quantity (roughly samples in our environment), the U-Net learns to compose, producing new stitched trajectories at test time. We provide evidence that this data scaling induces a smaller receptive field of the diffusion model in Appendix A.5.2. This confirms that the ability to generalize composition of sub-sequences is another property that diffusion models learn from data at scale, adding onto concept composition [53], consistency [68] and grammar [23].
6 Discussion
6.1 What Are the Keys to Composition?
While our work attempts explain why existing works in diffusion planning achieve composition, a reasonable question to ask is what engineering insights our findings have. While our results indicate that achieving whole-sequence compositionality through architecture choices may be under-explored, we emphasize that different choices are valid in different contexts. For instance, in environments where reactive control is very effective, generating long-horizon sequences may not be important, so frequent replanning with a short horizon may work better. In environments where data is abundant, using a higher-capacity model might be appropriate. We provide additional discussion of our findings in Appendix A.6.
6.2 Limitations and Future Work
Our work has several limitations. First, we only consider navigation environments with little complexity in the dynamics or observation space. While we motivate this in Section 5, evaluating our findings in a more realistic robotics environment would be useful future work. Second, our guidance strategy was limited to inpainting, which is effective in environments with clear endpoint goals, but less useful when goals are abstract (such as with language-conditioned agents) or multi-modal. Incorporating a way to leverage other guidance strategies like classifier-based [30] or free [31] guidance into our proposed high-compositionality diffusion planners is potentially fruitful future work. This would also let us fit our composition framework into the diffusion policy [13] setting, where composition of atomic skills may still be useful, but where a goal state cannot be inpainted because the diffusion model generates future actions and not future states.
Appendix A Appendix
Contents
A.1 More Related Work
Prior to diffusion planners, other generative models were common in imitation learning [29]. Most modern approaches are either based around autoregressive decoder-only transformers [58], such as the Decision Transformer [12], ACT [25], or diffusion models [29]. A more comprehensive survey on Diffusion Planners was prepared by [63].
We draw from a rich body of work in diffusion model generalization, which typically examine image-based models. Approaches include examining the effect of the diffusion process noise [65], architectural biases [36, 52], and representation learning [35] among others.
Another work that similarly analyzes architecture and inference strategies for diffusion planning is [46], which interestingly finds that the best diffusion backbone are transformers, which have no inductive biases towards locality and usually have positional attention added. However, their work is about general performance in offline RL and not about emphasizing compositionality, which as discussed may not be an important properly for strong performance on the D4RL benchmark due to a lack of required compositionality [26]. Analyzing how transformers differ from CNNs with respect to compositionality remains unexplored as future work.
When analyzing replanning, we apply a common form of it where replanning simply happens after some fixed number of execution steps. However, more sophisticated approaches to replanning exist which may have different effects on composition. For instance, [71] adaptively re-plans when the estimated likelihood of the current partially executed trajectory falls too low.
A.2 Additional Experiment Details
A.2.1 Unconditional Completion Rates
| U-Net | U-Net+Aug | Eq-Net+PosEnc | Eq-Net | Replan (16) | |
| Trials | |||||
| Completion |
Here we report completion rates in the unconditional environment. Table 1 shows the completion rates for each algorithm. As shown, every algorithm completes it close to of the time, with only the full-sequence models producing illegitimate trajectories a small portion (roughly ) of the time.
A.2.2 Data Scaling Experiment Details
For our data scaling experiments, we used a different environment, consisting of point navigation between points arranged in a decagon in euclidean space. Figure 8 shows a visualization. Each trajectory was tasked with travelling from point-to-point for a total of 5 visited points per trajectory. Data was collected with the same Gaussian noise as in the grid environment, and the trajectory length was also states long. We trained each of our models in this context for gradient steps, with hyperparameters otherwise the same as in Table 2.
We ran each model times to approximate the distribution of the model’s output. The metric we report in 7 is the percentage of generations that are not seen in the training set, where membership in the training set is determined by finding which key-points a trajectory visits in which order, and comparing them to that particular model’s dataset.
Note that this metric is somewhat biased towards the models that can compose but are trained on less data, as the space of compositionally legitimate but unseen in training trajectories is larger than for the model that already has most of the support of legitimate trajectories in its training set. However, our results show that the threshold is fairly binary, as the very low data models will not or will only very rarely produce a novel composition, but past a threshold the models will consistently produce novel trajectories a large percentage of the time.
A.2.3 Algorithm Hyperparameters
Here we report the common hyperparameter we ran our experiments with, unless otherwise specified. Table 2 shows the full values.
| Hyperparameter | Value |
| EMA Rate | 0.9999 |
| Diffusion Steps | 100 |
| Optimizer | AdamW [45] |
| Base LR | 2e-4 |
| Weight Decay | 1e-5 |
| LR Schedule | Cosine |
| Predict Noise or Clean | Clean |
| Sampling Temperature | 0.5 |
| Batch Size | 128 |
| Gradient Steps | 500,000 |
| Training Horizon Length | 512 |
| Convolution Padding | Repeat |
| Solver | DDPM[30] |
| SDE Form | Discrete[30] |
Instead of jointly generating states and actions as in [34], we use a separate inverse dynamics model (as in [1]). This removes the need to tune action versus state weighing in the diffusion objective, and generally performs better than joint prediction [46]. We execute this inverse dynamics model with feedback, executing the action at each time step by giving the model the current state in the plan and the current state from the environment , such that .
A.2.4 Experiment Computing Resources
All experiments for this work were done on a laptop with a 13th Gen Intel Core i9-13950HX and NVIDIA GeForce RTX 4080 Laptop GPU. The final models used to produce the results shown in the paper were trained over roughly 2 weeks, and model inference to generate the final results took roughly 1 week.
A.3 More Environment Details
A.3.1 Grid Environment Details
Our main environment is a small grid where a point agent navigates from one part of the grid to another while avoiding a series obstacles. The observation is the x-y coordinates of the agent, and the action is a x-y step which is normalized to a small step size. In the middle of the maze are several obstacles, which will kill the agent if it stays inside of the block for several steps in a row. Figure 1 shows examples of our environment, including points navigating to and from different locations in the grid. Navigating from one corner of the grid to the other requires steps out of a total allowed episode length of , challenging the diffusion planner’s long-horizon planning. The obstacles also ensure the trajectory must remain compositionally legitimate over the whole horizon, as accumulating error can cause the agent to wander off of the grid to its death.
A.3.2 Grid Environment Possible Trajectories
Lemma 1: For a grid of size where is the number of blocks along the grid’s height and width, the number of possible topologically distinct trajectories that are optimal in length is upper-bounded (inclusive) by .
Proof: First, observe that the number of optimal topologically distinct trajectories for any start-goal pair that is not from a corner to the opposite corner will be upper bounded by those that are. This is because they will have a smaller number of segments to traverse, meaning all of the analysis below applies with a smaller asymptotic quantity. This means finding the maximum number of distinct optimal trajectories in the corner to opposite corner case will upper bound the same quantity for any other start-goal pair.
Assume we are navigating from the top left to the bottom right corner for notational purposes. Observe that for any optimal path, the agent will move right times and down times. This means all possible trajectories can be represented by a binary string of length where there are an equal number of s and s. The size of this set is , because from positions we choose to be filled with either s or s.
A.3.3 Data Collection
In the unconditional setting, each trajectory starts in the top-left corner and navigates to the bottom-right corner, randomly switching from moving either downwards or right at each intersection. In total, our unconditional dataset includes unique trajectories out of a possible trajectories.999See Appendix A.3 for proof. This data quantity means our diffusion planner must generalize from a small number of samples to be able to generate samples in the (much larger) space of unseen but legitimate trajectories.
In the conditional setting, we replicate the above data collection process, but for all four corners, to ensure a diverse set of subsequences. We additionally add data starting from each edge intersection point going towards other randomly selected edge intersection points. This ensures that the dataset contains a wide range of subsequences, but that the majority of start-goal states are not seen during training. This dataset contains trajectories.
A.4 Further Architecture Details and Experiments
A.4.1 U-Net is Non-Local and Not Shift Equivariant
U-Net is Not Local: Here we support the claim made in the main text that the U-Net architecture proposed by [34] is not shift equivariant and displays heavy non-locality during the denoising process.
To show this, we follow the procedure of prior works in analyzing diffusion model local field receptiveness by calculating the gradient of an intermediate sample point with respect to the input sequence. We took our final trained U-Net and Eq-Net models in the unconditional environment and generated samples. For each of the denoising steps, we calculated the gradient of the middle pixel in the sequence with respect to the whole input sequence (possible because of auto-differentiation), and averaged these gradients across the generations for each time step. Formally, for a diffusion model where is the next de-noised sample, is the previous sample, and is the denoising step, we calculate for each diffusion time-step from . Because we generate a 1-D sequence, we can plot these gradient maps as an image, where the x-axis is the denoising step and the y-axis represents the sequence gradient.
Figure 9 shows these gradient maps. Visually, it is clear that the U-Net attends early in the denoising process to parts of the sequence far away from the centre pixel, even though the influence is much less strong than the pixels directly surrounding the centre. The size of the receptive field early in training spans far, containing non-zero information as far as the boundary states of the trajectory. On the other hand, Eq-Net has a hard boundary roughly pixels to each side of the centre pixel past which the model cannot attend to in a single denoising step. Like in prior works [36, 52], our diffusion models’ receptive fields progressively shrink during training, starting large and ending small.
We also plot a rough estimate of the local receptive field size at each denoising step. We do this the same way as [52]. We set a threshold value of the th percentile of the value of the absolute gradients in the image, find the window centred around the middle pixel that contains all values in that threshold for a given denoising step, and then plot the size of that window. Figure 10 shows the results of this, smoothed with a Savitzky-Golay filter. This plot shows that as expected, the local receptive field of the U-Net has a larger maximum value, approaching pixels wide at its max, while the Eq-Net is architecturally capped at .
This is a somewhat surprising result, given that [34] specifically highlights the local receptiveness from convolutions as supporting composition. Given that our results in both the conditional and unconditional setting found that the U-Net architecture struggles with composition, we directly analyzed the network during the diffusion process to gain some insight into whether the model is sufficiently local. We suspect part of the reason for this is because while stride 1 convolutions are locally receptive if applied to the whole sequence, pooling operations and long-stride convolutions both down-sample the image substantially, such that the internal trajectory length at the bottle-neck of the U-Net may be small enough that even a small convolutional kernel is capable of attending to information far across the sequence.
U-Net is Not Shift Equivariant
Strictly speaking, empirical evidence that U-Net is not shift equivariant or that Eq-Net is should not be required, as it directly follows from known properties of their architecture: the U-Net has down-sampling strides which break shift equivariance [70], and Eq-Net has only 1-length strides with no pooling layers which preserves convolutions’ shift equivariance.
For the sake of visual demonstration, we show this by taking our trained U-Net and Eq-Net models in the unconditional environment and giving the model an input sequence of all zeros, besides a patch of Gaussian noise in the middle of the trajectory. We then shift this Gaussian patch by pixels and analyze the output of each model. Formally, we take two initially empty trajectories and add a patch of Gaussian noise pixels wide to the middle of the first trajectory: , and add the same Gaussian noise shifted by pixels from the middle to the second trajectory: . Finally, we run each through a given diffusion model at : . A model with shift equivariance should find that the pixels of the output de-noised trajectory corresponding to the same patch as the Gaussian noise should be equal, that is, .
Figure 11 shows the output of each model at the original input position and the input position shifted to the right, as well as the difference between the outputs when the shift is corrected for. The U-Net architecture has substantial differences in the output when shifted, shown as a difference in the rightmost plot, while Eq-Net retains the same output but translated.
A.4.2 Engineering Details for Eq-Net
The main purpose of Eq-Net is not to propose a competitive architecture, but to show a proof of concept that locality and shift equivariance can be baked into architecture to enable stitching. Given this, we did not aggressively tune the architecture parameters of Eq-Net, instead finding values for network depth, kernel size, and channel size that worked well in our benchmark. However, we found it has some limitations and advantages that we will discuss here.
Eq-Net is substantially more memory efficient than the base U-Net model, with the base U-Net having million parameters and Eq-Net only having million parameters.
Despite this, Eq-Net was around half as efficient at both training and inference time. We suspect the former is because the channel sizes grew to be much larger in the intermediate representation layers of the U-Net ( at maximum) while Eq-Net has a fixed channel size of throughout. The latter is likely because Eq-Net is layers deep, which prevents efficient forwards passes because fewer operations can be parallelized.
Future work adapting these findings into a more practical architecture might investigate using adaptive polyphase sampling [6], anti-aliasing layers [70], or other modern tricks for incorporating shift equivariance into CNNs. Additionally, the depth of the network, way the layers of differing kernel sizes are layered, and channel size can likely be greatly optimized to increase computational efficiency and model capacity.
A.4.3 Kernel Size Tuning
The main parameter that requires tuning in Eq-Net is the kernel expansion rate (KER) which determines how fast the kernels grow deeper into the network. In Eq-Net, the kernels in each layer starts as small as possible (3) and grows 2 larger every KER layers, meaning for a fixed depth of , the maximum kernel size the network has is determined by . Thus, a small KER corresponds to a network with more larger kernels, and a larger KER with many small kernels.
Broadly, we found that tuning KER trades off between overall plan consistency (how likely the trajectories were to be locally consistent and stay outside the obstacles) and compositionality. To show this, we trained Eq-Net models in our unconditional setting with 3 KER values: 2, 5, and 12. Our results are shown in Table 3. We found that the intermediate value of 5 generated both a strong diversity of plans as well as reliably well-composed plans, while making the model too non-local made the model too local failed to consistently produce legitimate composed trajectories.
| Large (KER = 2) | Medium (KER = 5) | Small (KER = 12) | |
| Unique Trajectories | |||
| Completion Rate |
We also generated denoising gradient graphs of the small and large kernel Eq-Nets, like in our analysis of U-Net and our final Eq-Net architecture. Figure 13 shows the gradient graphs. As expected, Eq-Net with large kernels is attentive to information across the trajectory, and interestingly produces a gradient map extremely similar to the default U-Net, while Eq-Net with small kernels has an extremely limited receptive field.
Our experiments with the default U-Net architecture and our Eq-Net architectures in total confirm our earlier hypothesis that while CNNs are meant to have strong inductive biases to encourage composition, the particular set of decisions made by [34] means at small data amounts they can still overfit heavily and attend to non-local informatin,
A.4.4 Kernel Diversity
One thing that we found was important when designing Eq-Net was to include both very small (such as k=3) along with the larger kernel layers. When we used only medium sized kernels, it would lead to trajectories that were oddly "ungrounded" to the navigation plane, and would look correct in shape but would be shifted such that the resulting trajectory would nearly always hit obstacles. Figure 14 shows the visual results of this. We found that increasing the kernel size from small to large during training helped, somewhat mirroring the process in normal CNNs where fixed-size kernels stride over the whole image initially (capturing fine detail) and then smaller subsets after each downstride (capturing coarse detail). However, we did not explore whether the exact ordering of the kernel layers in Eq-Net matters or not.
A.4.5 Positional Encoding
When testing how our small receptive field model performs when it has access to positional information for each state, we use the same positional encodings as in [66], with slightly varied parameters. Figure 15 shows a visualization of how we add positional encoding to the Eq-Net architecture. These encodings turn the position of each state into waves forming a geometric progression from which the position of a particular state can be inferred. The encodings are added after the first residual block. Specifically, the positional encodings are defined as
We used a maximum length of instead of used by [66] because of the shorter length of our sequences.
It is worth noting that this strength of positional encoding (where the network is given enough information to precisely attend to the location of a state in the sequence) is likely quite a bit stronger than the positional attention from CNN aliasing. We included this as a representation of the worst-case scenario, to show the upper bound of models that are designed to attend to position, such as diffusion transformer architectures [56].
A.5 More Experiments
A.5.1 Overfitting Analysis
Here we analyze the effect of overfitting. A common strategy in all machine learning to increase generalization is to perform early stopping, where training is ended before loss on the training set plateaus. Here, we run all of the models shown in the unconditional experiments at gradient increments of and run each model for samples, tracking both their completion rates and number of unique trajectories. Figure 16 shows that overfitting indeed reduces compositionality in the small-data setting, as models early in training achieve both high completion rates (indicating the generated plans are kinematically feasible) as well as more novel trajectories than later in training. However, the models that show stronger compositionality (such as Eq-Net) retain this property late into training (i.e, the losses in compositionality from overfitting plateaus). This hopefully indicates that the proper inductive biases from architecture choices can avoid the overfitting problem.
A.5.2 Does Data Scaling Increase Locality?
In the Results section, we found that scaling data increased the compositionality of our base U-Net model, even though it does not have a strong inductive bias towards locality. To see if the compositionality induced by data scaling is attributable to data increasing locality, we analyzed the models as done previously in Appendix A.4.1 by recording the gradient of the middle pixel for samples across each denoising step for each model. We then plot these directly, and also plot the estimated size of the local receptive field for each model. Following [52], to plot the receptive field we set a threshold value of the th percentile of the value of the absolute gradients in the image, found the window centred around the middle pixel that contains all values in that threshold for a given denoising step, and then plotted the size of that window.
Figure 17 shows the gradient maps. Note that we name each data scale, with the amount of data corresponding to each name appearing in parentheses (SMM = Small Medium Medium, SM = Small Medium, SSM = Small Small Medium). Visually, it can be seen that the receptive field of the largest model (Full) is smaller than the field of the smallest model (Small) early in the denoising process. The small model attends to information clearly at the bounds of the image around step 20 of the denoising process, while the full data model does not have substantial gradient information past around 150 pixels away from the centre on either end.
Figure 18 shows the plotted local receptive fields, smoothed with a Savitzky-Golay filter. The model trained on the highest data scale has a clearly smaller receptive field compared to the models trained on less data. This indicates that the reason why data scaling may help composition is because it trains the model to have a smaller receptive field, highlighting locality’s importance for composition and the usefulness of our theory. More work should be done to understand why data scaling specifically induces such a property.
A.6 More Discussion
A.6.1 Broader Impacts
Our work does not have any broader societal impacts beyond the general impacts that sequential-decision making research may have on automation or enabling more intelligent agents. While our work is limited to planning, some of our finding might generalize to diffusion models more broadly, which have implications for copyright/artist’s rights, deepfake generation, et cetera.
A.6.2 How Should We Select a Method for Composition?
While we focus on the whole-sequence generation case, it is not clear what the strongest method is to generate heavily temporally correlated, sequential data. There is evidence that models trained only or mostly on auto-regressive prediction can still learn to do sophisticated forms of forwards and backwards planning [43]. Complex sequential generation models in video sometimes incorporate a form of auto-regressive prediction [62] but often generate the whole sequence in parallel with diffusion [60, 11].
Hierarchical diffusion appears to be a strong way to incorporate both the benefits of long-horizon planning and short-horizon consistency. Besides the empirical benefits shown in the hierarchical diffusion planning papers discussed in 2, more evidence exists that hierarchical models can learn to minimize dynamics error in fewer training steps [19]. One downside of these methods is that how the task is hierarchically decomposed either needs to be set manually, requiring more tuning, or must rely on a more complex algorithm [42] which increases implementation cost and other tuning. Image models have generally shown that forgoing manually set hierarchy and instead letting models jointly learn both high level structure and low level detail can work well given proper data and architecture decisions.
One consideration we do not explicitly examine is that using locality to achieve composition necessarily hurts the model’s ability to use important information from prior states. Long memory can be both a blessing and a curse, as it has found to increase a policy’s ability to handle very distant observations [51] but can also replicate undesired correlations (i.e, correlations that are circumstantial and which the model should not replicate) between past observations and future actions [16, 67]. The lack of compositionality of non-locally attentive models can be thought of as a form of causal confusion, as they "confuse" specific transitions as being associated with state visitations in the distant pass that may not hold relevance for the future plan. However, recent work has found that taking a linear combination of the conditioned diffusion steps on a variety of past history lengths may preserve the benefits of both long-horizon planning and short-horizon composition, so it may be possible to combine the benefits of both [62].
A.6.3 Explaining Composition in Other Diffusion Planning Works
Here, we apply our framework to a couple of other works in the diffusion planning literature.
Hierarchical Diffusion Planning: As discussed in Section 2, several works approach the composition problem by separately generating sub-sequences with a diffusion model, and then composing them. Exactly how these low-level subsequences are "orchestrated" into a single coherent plan varies, but each exhibits aspects of locality and shift equivariance.
[10] propose Hierarchical Diffuser, which uses two separate diffusion models: a high level one which generates a "plan sketch", and a low level one which generates a sequence connecting the states in the plan sketch to "fill-in" the high-level plan. Functionally similar was the prior HDI paper from [42], which has a different form of generating sub-goals unrelated to diffusion. These approaches enable a form of shift equivariance by data augmentation: because the low-level subsequences generated by the low-level diffusion model are randomly cut out of the longer trajectories in the training data, they have their relative position within longer sequences removed. This also enables locality by disentangling the joint prediction of states within a particular subsequence to any other states outside of it, so each subsequence functions as a small local window without connectivity to previous or future states, preventing confounding from preventing stitching. [10] also confirms some of our findings that Diffuser with a larger kernel size often produces more coherent long-term plans (4.3) but that this trades-off with worse ODD compositional generalization (Appendix E).
[48] proposes a hierarchical method called CompDiffuser which, like the Hierarchical Diffuser paper, generates sub-sequences. Instead of having a separate high-level planner, they have each sub-sequence condition its generation on its immediate surrounding neighbour subsequence (3.2). This preserves shift equivariance in the same manner as Hierarchical Diffuser, but maintains locality through this process of having each sub-sequence only attend to its neighbours at each step, which stops causal confusion.
Diffusion Forcing: Some recent works extend diffusion into very long-horizon generation through using a combination of auto-regressive sampling and diffusion sampling called diffusion forcing. Diffusion forcing denoises over a whole sequence, but iteratively denoises near-future states more aggressively, leading to the sequence being progressively denoised from start to finish instead of the entire sequence being denoised consecutively. As discussed in the main text, short-memory and future prediction enforces locality by reducing the ability for long-horizon confounding effects to prevent composition.
Results from [9] shows that this approach can enable compositionality, but only when keeping short or no memory and only generating short plans, essentially using the diffusion forcing model to approximate short-horizon Model Predictive Control (E.2). When extending the generation length or memory, they find that it instead reproduces existing trajectories.
Follow up work from [62] extends Diffusion Forcing by replacing their RNN-based backbone with a transformer, and introducing the idea of combining the scores of the same model conditioned with several history lengths to try to achieve both long-sequence compositionality and short-term reactivity. They tested this approach on a task where a robot must take a fruit in a slot and move it to another slot. This task requires both memory (of which fruit the model should move) and short term reactivity (a human perturbs the robot mid-trajectory). However, the training data does not have both memory of the original fruit location and examples of recovery behaviour from the human perturbance, requiring composition of training sub-sequences. They found that the model that keeps the full trajectory history overfits to the training datasets showing the original fruit, while only using short-horizon models would compose recovery behaviour but forget the original task. Ultimately they found combining the score produced trajectories that both remembered important fast information and were capable of composing sub-trajectories for reactive control. This supports our argument highlighting the importance of locality for composition, but highlights that locality alone can run into issues in long-horizon tasks, as discussed in Section 6.1. For details, see Sections 6.4: Task 3, C.8, and D.4.
Diffusion Policies/VLAs
Perhaps the most popular diffusion-based sequential decision making algorithms today are various forms of Diffusion Policies [13]. These are essentially similar to the prior Diffusion Behaviour Cloning framework [55] where the diffusion model acts as an explicit policy and generates an action, except that Diffusion Policy predicts a short horizon of actions that are executed. This framework of predicting a short horizon of actions with diffusion is also used by most modern foundation-model Vision Language Action models like Pi0.5 [32] or Gr00T [4].
This differs from the diffusion planning framework in that a long-horizon plan is not explicitly generated, so any long-horizon task guidance is obtained from conditioning the policy (like in goal-conditioned BC) rather than from explicit planning. Thus, in these models compositionality essentially comes from the same place that it does in TD learning: because no planning is done, the model composes through replanning. This replanning, like any explicit one-step or few-step policy, assumes the environment is Markovian and so does not include temporal information or preserve long history as conditioning, thus preserving shift equivariance and locality as we discussed in 3.2.
There is some limited early evidence that when trained from scratch on very small datasets, Diffusion Policies may perform learning more similar to memorization with nearest-neighbour interpolation than "generalization" in any meaningful sense, and thus obtain strong apparent trajectory composition just through being very reactive [27]. However, these results have only been found in very small data settings, so it is likely that they do not apply to larger VLAs or to multi-task Diffusion Policies trained on a heterogeneous mix of tasks.

### v3

@@
-What Do You Need for Diverse Trajectory Composition in Diffusion Planning?
+What Do You Need for Compositional Generalization in Diffusion Planning?
 Abstract
-In planning, stitching is an ability of algorithms to piece together sub-trajectories of data they are trained on to generate new and diverse behaviours.
-While stitching has historically been a strength of offline reinforcement learning, recent generative behavioural cloning (BC) methods have also shown proficiency at stitching.
-However, the main factors behind this are poorly understood, hindering the development of new algorithms that can reliably stitch.
-Focusing on diffusion planners trained via imitation, we find two properties are key contributors for composition: shift equivariance and local receptiveness.
-We use these two properties to explain architecture, data, and inference choices in existing generative BC methods based on diffusion planning, including replanning frequency, data augmentation, and data scaling.
-Experimental comparisons show that (1) while locality is more important than shift equivariance in creating a diffusion planner capable of composition, both are crucial (2) enabling these properties through relatively simple architecture choices can be competitive with more computationally expensive methods such as replanning or scaling data, and (3) simple inpainting-based guidance can guide architecturally compositional models to enable generalization in goal-conditioned settings.
-1 Introduction
-The recent resurgence in generative diffusion models [58, 30] has inspired a new wave of work investigating the potential for these models to act as planners for robotic agents [34, 13].
-One surprising ability these diffusion planning models have is compositionality: the ability to take component elements seen in samples of their dataset and compose them in new, kinematically feasible ways [20, 44].
-One form of compositionality especially important in diffusion planning is stitching, where some diffusion planning methods are capable of composing sub-trajectories into new trajectories distinct from the training data.
-Historically, stitching has been seen as one of the primary benefits of temporal difference (TD) learning [38], both in offline and online RL.
-Planning with diffusion models thus potentially offers some of the stitching benefits of TD learning without needing to handle RL-specific challenges.
-Despite their recent success in vision, when and why diffusion models compose in the planning domain is not well understood.
-Contrary to toy experiments in prior works [1, 34], the convolutional neural network (CNN)-based backbones commonly used in diffusion planners often struggle with composition [9], instead regressing to memorization of training examples [2].
-This gap between the promise of diffusion planners to enable stitching and reality that they frequently struggle with composition is crucial to address, as diffusion offers an appealing alternative to other stitching approaches.
-RL suffers from many challenges when stitching such as the deadly triad [33, 22, 64] and slow temporal credit assignment [12], while non-generative BC stitching approaches require nettlesome data augmentation [26].
-The capability to sample a diverse set of new trajectories at test time is important property for planners broadly.
-Reasons for this include being able to select solutions from a candidate set [69], encouraging multi-modal behaviour [40], or generating diverse data to train imitation learning or offline RL algorithms [38].
+In policy learning, stitching and compositional generalization refer to the extent to which the policy is able to piece together sub-trajectories of data it is trained on to generate new and diverse behaviours.
+While stitching has been identified as a significant strength of offline reinforcement learning, recent generative behavioural cloning (BC) methods have also shown proficiency at stitching.
+However, the main factors behind this are poorly understood, hindering the development of new algorithms that can reliably stitch by design.
+Focusing on diffusion planners trained via generative behavioural cloning, and without resorting to dynamic programming or TD-learning, we find three properties are key enablers for composition: shift equivariance, local receptive fields, and inference choices.
+We use these properties to explain architecture, data, and inference choices in existing generative BC methods based on diffusion planning including replanning frequency, data augmentation, and data scaling.
+Our experiments show that while local receptive fields are more important than shift equivariance in creating a diffusion planner capable of composition, both are crucial.
+Using findings from our experiments, we develop a new architecture for diffusion planners called Eq-Net, that is simple, produces diverse trajectories competitive with more computationally expensive methods such as replanning or scaling data, and can be guided to enable generalization in goal-conditioned settings.
+We show that Eq-Net exhibits significant compositional generalization in a variety of navigation and manipulation tasks designed to test planning diversity.
+I Introduction
+The advent of diffusion models [36] has inspired a wave of work using these models as planners and policies in robotic systems [40, 13].
+Image and video diffusion models exhibit strong compositionality: the ability to take component elements seen in samples of their dataset and compose them in new, feasible ways [37, 21, 52].
+When and why diffusion models for planning and control compose, however, is not well understood.
+Diffusion Policies [13], which generate short action chunks, naturally compose behaviours through frequent replanning.
+However, Diffusion Planners [40], which generate long-horizon plans, often struggle with composition [28, 9], instead reproducing exact training trajectories.
+Composition in diffusion planners is crucial to address, as the capability to compose behaviours into new trajectories at test time is important property for planners broadly.
+Reasons include allowing for better candidate set selection [79], encouraging multi-modal behaviour [47], and generating better data for offline reinforcement learning [45].
 To address this gap, we analyze a variety of approaches that induce composition in diffusion planners.
-We find that there are two key ingredients: local receptiveness [52] during generation, and shift equivariance of states in a trajectory [36].
-Through experiments in navigation environments, we show that although most current diffusion planning models do not possess this compositional capacity natively, it can be induced through architecture changes, inference choices, data augmentation strategies, or data scaling without needing dynamic programming or TD-learning.
-We further show that the diversity of trajectories this compositionality enables can be selectively sampled from using inpainting, allowing the use of diffusion planners only trained on imitation as goal-conditioned planners, similar to offline goal-conditioned RL.
-Our primary contributions are as follows:
-Showing that local attention and shift equivariance, previously shown to mechanistically explain creativity and generalization in image diffusion models [36], also function as ingredients for composition in diffusion planners.
-However, unlike the image diffusion case, they are not the only factors: the inference method used to sample trajectories from diffusion planners also plays a significant role.
-Demonstrating that when induced through architecture changes, the stitching capabilities of whole-sequence diffusion can be effectively guided through simple inpainting to achieve strong goal-conditioned performance on unseen start-goal pairs.
-Our work provides more clarity behind what components are necessary to guide diffusion planners to compose trajectories at test-time successfully.
-We hope this will inspire future algorithmic and architectural developments in compositional diffusion planners.
-2 Related Work
-Diffusion Models: Diffusion models [61, 30] offer an efficient way train models that can sample from complex distributions.
-Diffusion models train from data points drawn from some target distribution by taking a sample and iteratively adding Gaussian noise, such that , called the forward diffusion process.
-The diffusion model learns the reverse diffusion process, where some model (usually a neural network) parameterized by is trained to learn by minimizing the MSE loss between the model’s output and the reverse diffusion step: .
-Sampling from the new distribution is performed by first sampling from the Gaussian and using the model to reverse-sample the diffusion process: [50, 7].
-Diffusion Planning: Introduced by [34], diffusion planners model trajectories with diffusion.
-A diffusion model predicts an upcoming state or state-action trajectory horizon time-steps long, conditioned on a starting state or state memory of length , in an environment with max steps.
-From the generated plan, actions are executed for a fixed horizon before generating a new trajectory.
-In this work, we are mostly interested in the whole sequence planning domain, where , and no replanning occurs.
-Diffusion planning should not be confused with the diffusion policy [13] formulation where the diffusion model generates a short sequence of actions with high reactivity from frequent re-generation of actions.
-We elaborate on this distinction further in Appendix A.6.3.
-Stitching in Imitation Learning: While stitching is often viewed as a strength of TD learning [38], other work analyzes stitching in imitation learning [8, 28].
-We were inspired by [26], which analyzes stitching through a compositional generalization perspective.
+We find that there are two key ingredients: the ability to use sub-skills at any point in a sequence (shift equivariance [42]), and a locality bias towards attending to nearby states when generating a state in a trajectory (local receptive fields [62]).
+We show that previous methods achieve composition through these two ingredients, identifying composition can be ensured through careful inference choices, data augmentation strategies, or data scaling.
+Using these two ingredients, we develop a neural network architecture for the diffusion planning denoising network called Eq-Net, which is simple, produces diverse trajectories, and can be guided to enable strong goal-conditioning.
+We show Eq-Net produces more diverse trajectories than other architectures in a variety of navigation and robotic manipulation settings.
+Our contributions are as follows:
+We show that local receptive fields and shift equivariance are key enablers for composition in diffusion planners, as has been shown in the image generation domain [62, 42].
+However, unlike the image generation case, they are not the only important factors: the inference method used to sample trajectories from diffusion planners also plays a crucial role in achieving compositional generalization.
+We show that stitching achieved through data scaling, as opposed to architectural choices, still induces local receptive fields in models.
+We demonstrate that when induced through architecture changes, the stitching capabilities of whole-sequence diffusion can be effectively guided through inpainting to achieve strong goal-conditioned performance on unseen tasks.
+Our work aims to provide clarity behind the key factors that guide diffusion planners to compose trajectories at test-time, informing future algorithmic and architectural developments in compositional diffusion planners.
+II Preliminaries and Related Work
+Diffusion Models: Diffusion models [71, 36] are generative models trained on data points drawn from some target distribution by taking a sample and iteratively adding Gaussian noise, such that , called the forward diffusion process.
+The diffusion model learns the reverse diffusion process, where a neural network is trained to learn by minimizing the MSE loss between the model’s output and the reverse diffusion step: .
+Sampling is performed by sampling from the Gaussian and using the model to reverse-sample the diffusion process by taking repeated steps of size from : [60, 7].
+Many different formulations to diffusion model prediction exist, but most works either predict a clean sample from noise [49] or the residual noise to remove [36].
+Diffusion Planning: Introduced by Janner et al.
+[40], diffusion planners model entire trajectories by sampling from diffusion models.
+A diffusion planner generates an upcoming state trajectory consisting of a sequence of states horizon time-steps long, sometimes conditioned on a state memory of length in an environment with max steps.
+Actions from the plan are executed for a horizon before generating a new trajectory.
+Here we mostly investigate the whole-sequence planning domain, where , and no replanning occurs.
+The more common Diffusion Policy [13] formulation, where the diffusion model generates a short sequence of actions with high reactivity from frequent re-generation of actions, can be thought of as a special form of diffusion planning where .
+We elaborate on this distinction further in Appendix VI-H3.
+Compositionality in Diffusion Planners: Many approaches exist to explicitly induce compositionality in diffusion planners.
+Hierarchical approaches use diffusion to stitch existing dataset trajectories [48], generate subsequences jointly conditioned on each-other [57], or separately generate high and low level plans[10, 50].
+By contrast, we target whole-sequence compositionality, where composition does not occur from explicit stitching but from implicit model behaviour, similar to image diffusion.
+Benefits of a monolithic planning approach includes reducing algorithmic complexity, sidestepping potentially harmful planning hierarchy [2], and enhanced performance in long-horizon tasks [9, 54].
+Generalization in Image Diffusion Models: Work in Diffusion model generalization typically examines image-based models.
+Approaches include examining the effect of the diffusion process noise [74], architectural biases [42, 62], data statistics [56] and representation learning [41] among others.
+We extend these works by connecting their observations to robotics and showing that through introducing architectural biases, we can enable compositional models in low-data regimes.
+III Methodology
+III-A Do Existing Diffusion Planners Compose by Default?
+To motivate our problem, we observe that when standard diffusion planners are naively applied to a variety of tasks, the planners tend to exactly reproduce training trajectories.
+Figure 2 shows the rate that the two most common diffusion planner backbones, U-Net [40] and DiT [19], produce memorized trajectories in our testing environments designed to test compositionality, as we define in Section III-E.
+This is surprising, given that diffusion models exhibit creativity in image and video domains [30].
+However, diffusion planners are often trained and evaluated differently than image diffusion models.
+Image diffusion models are trained to produce whole images at once.
+Meanwhile, in diffusion planning people employ different inference and training techniques, usually introducing training augmentations or generating smaller subsequences.
+We outline the most common of these techniques in Figure 3.
+III-B Local Receptive Fields and Shift Equivariance in Score-Based Diffusion Models
+To unify prior explanations for compositionality in diffusion planners, we rely on two properties previously described for diffusion models in vision [42, 62].
+The first is local receptive fields, meaning the diffusion model’s denoising prediction for step in the trajectory is conditionally independent of far-away steps in the trajectory.
+Local receptive fields ensure that at each step, the diffusion model only attends to a small patch around part of a sample when denoising.
+However, through multiple denoising steps, information can propagate across a generated sample, shown in Appendix VI-B1.
+Definition 1: Locally receptive fields.
+Assume we are interested in generating a trajectory , and have a generative trajectory model whose output estimates the denoising function for individual states where indexes states .
+We say that has a field locally receptive to a state neighbourhood if for all trajectories and state positions within those trajectories , .
+The second is shift equivariance, which requires that the model ignore the absolute position of a state in the sequence as a condition during the denoising process.
+Definition 2: Shift equivariance.
+A trajectory score model is shift equivariant [59] if for all shifts on trajectories which uniformly change the position of states within the trajectory, .
+For diffusion planners, this can also be thought of as the diffusion model having consistent denoising predictions for similar sub-trajectory skills, regardless of their absolute position in trajectory.
+For instance, a diffusion planner should be able to de-noise a "pick" sub-skill for a robotic arm near the beginning or near the end of a trajectory, regardless of where that pick appeared in training trajectories.
+III-C Additional Factors for Composition in Diffusion Planners
+Our main insight is that these notions of local receptive fields and shift equivariance are similar to other strategies used to increase composition in diffusion planners.
+These other strategies are a crucial component to how modern diffusion planners compose [13, 57].
+Increasing Compositionality by Replanning at Inference: The first strategy we address is replanning.
+A trajectory planning model replans when conditioned on past observations it generates a trajectory of length , executes some action horizon , and then generates a new trajectory starting at the current state.
+Increasing extends the model’s effective planning horizon.
+Formally, given a generative trajectory model that generates trajectories of length conditioned on a history of length , each state in the trajectory will only be directly attentive to a local field of size , i.e, .
+Short-horizon diffusion planners, such as diffusion policies [13] achieve composition through this mechanism.
+However, this does not elucidate how longer-horizon planners compose.
+Note that fixed-memory replanning and local receptive fields in the reverse-diffusion process are not identical, as the many steps in the reverse diffusion process lets influence spread along a trajectory through many local denoising steps, shown in Appendix VI-B1.
+However, the similar mechanisms highlight how increasing the influence of nearby states in a given trajectory is important for composition.
+Replanning also ensures a form of shift equivariance, as the diffusion model must be capable of generating chunks at any time-step in the trajectory.
+Increasing Compositionality by Data Augmentation: Another strategy to increase compositionality is data augmentation, where data is modified during training to teach the model equivariance to a transformation [3, 69].
+The most common augmentation to encourage stitching is stitching together sub-trajectories in the training set randomly [27, 8, 34, 46].
+However, this strategy requires tuning the distance metric that determines when segments are close, and is infeasible for datasets with a large combinatorial set of legal stitches because such sets often grow exponentially.
+We instead focus on augmenting the data to provide shift equivariance.
+Specifically, we randomize the position of states in a trajectory during training to prevent the model from overfitting to a state’s position.
+We sample a random state in a trajectory, removing states in the past, and then pad the end of the trajectory with the final state, so that the trajectory is randomly "shifted" by a number of states.
+Formally, given a training trajectory with states , we sample a position from the uniform distribution .
+To create our new augmented trajectory , we shift the states in by positions to the left, such that for state indexes , .
+For the rest of the states , we pad the states with the final state: .
+This modified trajectory is then used in training.
+Increasing Compositionality via Architecture Design: The next strategy we address are architecture modifications.
+While equivariance can be learned through data augmentation, another strategy is to build equivariance to certain transformations through architecture changes[14, 15].
+The more modern DiT backbone is Transformer based [75], which does not have an inductive bias for local receptive fields when processing input sequences and has explicit positional encoding.
+While CNNs are often said to have locality and shift equivariance [40, 42], we found that the CNN based U-Net [40] has a globally receptive field and is not equivariant to shifts, shown in Appendix VI-E1.
+In fact, down-sampling and pooling operators common in CNNs provably break shift equivariance [80, 6].
+To analyze the effect that local receptive fields and shift equivariance have on composition, we propose a network called Eq-Net, a residual CNN [33], with two important design decisions:
+Convolutional layers without pooling or down-sampling guarantee equivariance to shifts [80].
+Small convolutional kernels guarantee local receptive fields.
+See Appendix VI-E2 for further elaboration and an architecture diagram.
+Increasing Compositionality via Scaling Training Data: The last strategy we address is scaling training data, which increases diffusion model generalization [78] and compositionality [63].
+This increase in compositionality does not come from memorizing more of the sample space, but from learning the rules governing composition in a particular area [25, 56].
+Our work is specific to robotics, where data is often sparse.
+Vision datasets range from samples [44] to millions [17, 77] of images or more.
+By contrast, robotics datasets are smaller, varying from a couple dozen trajectories [13] to a few hundred [58] or thousand [64].
+Given that most robotics datasets have relatively few samples, considering other approaches to generalization becomes more important than relying on data scaling.
+III-D Environment Descriptions
+We examine diffusion composition in a maze navigation task and robotic manipulation tasks.
+Our environments are designed to measure test-time stitching, with specific combinations of subskills able to be composed from the training data, but not seen in the same trajectory.
+Pictures of our environments are shown in Figure 4.
+Maze is a 2D point-maze navigation task where the point agent travels from the top left to the bottom right corner, taking different routes around obstacles.
+Combining sub-routes into a new path requires composition.
+Colliding with an obstacle kills the agent, requiring precise trajectories.
+Figure 4 shows a maze for clarity, but we used a larger maze in most experiments.
+Didactic is a manipulation task where a robot arm manipulates cubes.
+The left side of Figure 1 shows the training three trajectories of pick and places in two separate areas, and a trajectory connecting the two.
+Picking the cube from one area and placing it into the other requires composition.
+In Well-Plate, an arm must relocate blocks from one plate to another.
+Placing blocks into novel configurations requires subskill composition, both to compose pick and place sub-skills for individuals blocks together, but also the pick skill for a particular block with the place skill of another.
+In Block-Stack, an arm stacks six cubes.
+Not all stacking orders are seen during training, so stacking the cubes in new orders requires composition.
+In Lights, an arm presses buttons on a table at random to light them up.
+Novel trajectories require new permutations of active lights.
+We implemented the Maze task in Numpy [31], the Well-Plate Sim and Block-Stack tasks in RoboSuite [82], and the Didactic and Well-Plate Real tasks on a real Franka Emika Robot [29, 22].
+We adapt the Lights task from OGBench [64] with modifications.
+Data Collection for Diffusion Model Training: For each task besides the Didactic scene we collected between and trajectories by randomly sampling subgoal permutations and executing the corresponding trajectory through task-specific controllers.
+Each dataset only covers a small space of the potential legal task executions, allowing for novel generations from the diffusion planner, elaborated on in Appendix VI-D.
+III-E Measuring Trajectory Diversity and Novelty
+To quantify the diversity of each diffusion planner’s generations, we construct a general quantitative measure of the novelty of a trajectory.
+Our metric uses the structure of the task and training data to determine how novel a given trajectory is.
+Assume that each environment has a state represented by a set of primitives .
+Each trajectory modifies the state of the environment such that the final state can be represented by a set of these primitives.
+We manually design our environments around primitives that are easy to extract from the environment post execution of the trajectory.
+For instance, in our Maze task the primitives correspond to individual intersections in the maze.
+Each unique path through the maze is represented by the set of intersections it visits.
+Appendix VI-D3 elaborates on the primitive set for each environment.
+Applying this to each training trajectory means the training set can be represented by a list of primitive sets .
+When we generate a trajectory with a diffusion planner and execute it in the environment, it produces a new primitive representation .
+If s.t.
+, then we say that is memorized.
+This is because the generated trajectory is entirely composed of sub-trajectories seen together in a trajectory previously, and so no composition had to be performed.
+Otherwise, it is novel because producing that trajectory required composing elements from two or more training trajectories together.
+If (meaning the evaluation trajectory has the same number of primitives as the training trajectories), then we say that is complete.
+This is because the generated trajectory performed a number of subtasks equal to the ones seen in each training trajectory.
+Otherwise, it is partial because it only completed a smaller number of subtasks.
+Because we fix the number of subtasks each training trajectory performs, this lets us easily tell if a generated trajectory performs less subtasks than desired.
+The ideal compositional diffusion planner should produce trajectories that are both novel and complete.
+III-F Algorithm Setup
+For our experiments we use a 1D U-Net [40] and DiT [19] diffusion planners [40] as the base models.
+We use CleanDiffuser’s [18] implementation and hyperparameters.
+During inference, we predict state trajectories[1] and use a Cartesian operational space controller [43] to follow the trajectory.
+Appendix VI-C2 contains more details.
+IV Experiments
+Our hypothesis is that local receptive fields and shift equivariance are key ingredients in explaining compositionality in diffusion planning.
+We expect that different ways of incorporating each (training augmentations, replanning, architecture choices, etc.) should increase compositionality.
+However, it is not known which of the two properties is more important, whether both are needed, or which way of inducting the two properties increases compositionality the most effectively.
+It is also not known how scaling data is related to local receptive fields.
+Our experiments aim to address these questions.
+IV-A How Does Architecture Affect Composition?
+Our primary question is if the lack of local receptive fields and shift equivariance in diffusion planner architectures hurts compositionality.
+We trained U-Net, DiT, and Eq-Net on all of our environments, letting us directly compare the effects of architectural biases on composition.
+We ran the resulting planners between and times in each environment, and calculated the diversity and completeness of each trajectory.
+Figure 5 shows the results on five of our six environments111Due to Didactic only being completion-based, we defer results to Table I., where we report the percentage of model samples that exhibit completeness (completing the same number of subtasks as in each training sample) and novelty (whether samples must have combined elements from different training data samples).
+Eq-Net shows increased novelty in all environments tested, confirming our hypothesis on the importance of local receptive fields and shift equivariance.
+IV-B Locality vs Shift Equivariance: How Much Do They Matter?
+Eq-Net is designed to have local receptive fields and be shift equivariant, meaning it is not clear which of the two properties is most important.
+We trained additional models in a version of our Maze environment, one with only local receptive fields and the other with only shift equivariance.
+Specifically, we train a non-local model to have approximate shift equivariance through training augmentation (U-Net+Aug, described in Fig 3) and a local model without shift equivariance by adding sinusoidal positional embeddings [75] (Eq-Net+PosEnc, described in Appendix VI-E5).
+Figure 6 shows how the diversity of each model’s outputs increases as more samples are collected.
+Notably, Eq-Net+PosEnc performs better than U-Net+Aug, indicating that local receptive fields are a comparatively more important factor than shift equivariance.
+However, the combination of both (Eq-Net) achieves the best performance, indicating that both ingredients are crucial.
+We show additional experiments that increasing Eq-Net’s kernel size decreases compositionality in Appendix VI-E3, validating our theory that receptive local fields are important for enhanced composition.
+IV-C Replanning During Inference vs Architectural Design
+As discussed in Section III-C, compositionality from generating smaller chunks of a trajectory, either through replanning or a hierarchical planning method, is a common way to enhance compositionality in diffusion planners.
+Our theory predicts these algorithms will outperform monolithic planning, as their inference and training methods enforce local receptive fields and positional equivariance.
+We compare Eq-Net to visual Diffusion Policy [13], which generates short-horizon chunks and frequently replans, and CompDiffuser [57], which generates several chunks conditioned on their neighbours which are then stiched into a longer plan.
+We trained both algorithms with their original hyperparameters on our Robosuite and Navigation tasks.
+Results are shown in Figure 7.
+As expected, these chunking methods compose well, performing comparably to Eq-Net on our manipulation tasks.
+CompDiffuser struggles on the Navigation environment, as the chunking strategy often led to sub-par trajectory quality along the overlap points between trajectories which would guide the point agent into obstacles.
+This shows that monolithic planners can preserve overall trajectory fidelity that chunking methods struggle with.
+IV-D Compositionality via Data Scaling vs via Architecture
+To confirm that data scaling can increase compositional generalization, we produced larger and smaller versions of our Maze dataset.
+As expected, we found that simply increasing the amount of training data was often sufficient to increase a model’s ability to compose.
+Figure 8 shows that all models learn to compose with enough training data.
+However, Eq-Net’s built-in architectural biases help it generalize faster than the models without inductive biases.
+To examine how scaling training data relates to local receptive fields, we examined the receptive field of our models trained at different data scales.
+Following prior work [62, 42], we calculate the gradient of an intermediate sample point with respect to the input sequence.
+Formally, for a diffusion planner where is the next de-noised sample is the previous sample, the samples are length and state are indexed by , and is the denoising step, we calculate for each diffusion time-step from .
+We compute these gradients through the diffusion model exactly using auto-differentiation.
+Because diffusion planners generate a 1-D sequence, we can plot these gradient maps as an image, where the x-axis is the denoising step and the y-axis represents the sequence gradient from the start to the end of the trajectory.
+Figure 9 shows the gradient maps for each architecture trained on a small amount of data ( samples) and a large amount of data ( samples).
+To gain a lower-variance estimate, we average this gradient over samples.
+These show that the receptive field of the models trained on less data are larger for the non-local models (U-Net and DiT) than for the local Eq-Net, but as the amount of training data is increased the receptive fields look more similar and local.
+In particular, the receptive field starts out large early in the reverse diffusion process, but shrinks later in generation, with more shrinkage occurring for models trained on more data.
+We present the gradient maps for all models in Appendix VI-G1.
+To create a single quantitative measure of how local each model is, we follow prior works [62] and estimate the size of the receptive field by measuring the size of the centred window required to capture of the gradient magnitude.
+To estimate the total non-local behaviour of each model, we take the integral over the receptive field size across the entire trajectory generation.
+Figure 10 shows this measure of the strength a model’s local receptive fields for the three architectures at different data scales.
+There is a clear trend where increasing the amount of training data increases a model’s local receptive fields.
+Eq-Net also shows stronger local receptive fields across the board than U-Net and DiT, confirming the role that Eq-Net’s smaller kernels have in enabling stronger compositionality.
+Taken together, our findings show that even when a diffusion planner does not have inductive biases towards local receptive fields, scaling data will teach the model to have local receptive fields.
+Combined with recent findings that diffusion models learn different biases from different datasets [56], this suggests Eq-Net takes advantage of the fact that trajectory data is statistically more locally biased than globally biased.
+This is coherent with the fact that the rough Markovian property of trajectory data means individual states more strongly correlate with nearby states, and that in mixed behaviour/multi-task datasets long-horizon state correlations are weak.
+IV-E Are More Compositional Models More Steerable?
+One of the motivations behind planners with compositional generalization is that creating more diverse plans may be important when planning to solve a specific goal.
+While many guidance approaches exist [54], most works in diffusion planning use some form of conditioning to steer the model towards plans that solve a particular task [40, 1].
+To understand if our more diverse models are more guidable, we examined guidance in a variant of our Maze environment (Appendix VI-D4) and in a didactic manipulation environment (Section III-D).
+For both tasks we guide the models with inpainting like in prior diffusion works [55, 40, 71], where the starting state and desired ending or intermediate states are inpainted to the generated trajectory and fixed during denoising.
+Results for both goal-conditioned tasks are shown in Table I.
+Eq-Net far outperforms all other models in Maze, achieving an average completion rate, roughly higher than the model with only local receptive fields but not shift equivariance.
+To visualize the difference in composition for each model, we plotted the unique successful runs from random generations from each algorithm on one of the held-out start-goal pairs in Figure 11.
+Similarly, Eq-Net was the only architecture capable of producing a successful trajectory in our didactic environment, one of which is shown in the right-hand side of Figure 1.
+[…]
+| Model | U-Net | U-Net/Aug | Eq-Net/PosEnc | Eq-Net |
+| Completion | 0±0 | 0.44±0.0 | 13.39±5.0 | 83.8±6.4 |
+| Unique Paths | 0±0 | 0.09±0.0 | 11.16±1.0 | 116.8±4.8 |
+[…]
+| Model | U-Net | DiT | Eq-Net | |
+| Completion | 0 | 0 | 77 |
+In total, our goal-conditioned tasks show that models with compositional generalization are indeed more steerable, giving our findings use in guided planning methods.
+IV-F Remaining Factors for Diffusion Planner Composition
+While we cannot rule out other potential influences on compositionality, we found that some other phenomena had a small effect. In particular, we found
+Early stopping helps compositionality somewhat as seen elsewhere [5], but model bias matters much more.
+The number of denoising steps greatly affects trajectory completion rates, but not compositionality.
+Other algorithmic and parameter choices, such as using the discrete or continuous stochastic differential equation formulation, the diffusion sampling temperature, predicting noise versus clean samples.
+etc.
+did not make a difference.
+We defer further experimental results to Appendix VI-F.
+V Discussion
+V-A Explaining Composition in Other Diffusion Planners
+Our goal in examining the factors that lead to composition is to help explain why prior works compose (or fail to).
+We include a thorough discussion of other works in Appendix VI-H3.
+We find the majority of existing works achieve local receptive fields and shift equivariance either through frequent replanning [72, 9, 13] or some form of hierarchical decomposition where sub-trajectories are generated and then composed manually [10, 57, 50].
+This suggests that the form of composition seen in image diffusion models is mechanistically different from how most diffusion planners achieve composition, as most image diffusion models do not manually stitch together patches/frames.
+We hope our work proves the viability of a simple, purely architectural approach.
+We provide additional discussion and recommendations in Appendix VI-H.
+V-B Conclusions, Limitations, and Future Work
+Our work has several limitations.
+First, we only consider environments carefully designed to require stitching.
+While we motivate this in Section III-D, evaluating our findings in multi-task or even multi-embodiment robotics dataset would be useful future work.
+Second, our guidance strategy was limited to inpainting, which is effective in environments with clear endpoint goals, but less useful when goals are abstract (such as with language-conditioned agents) or multi-modal.
+Incorporating a way to leverage other guidance strategies like classifier-based [36] or free [37] guidance into our proposed high-compositionality diffusion planners is potentially fruitful future work.
+Acknowledgements
+We thank the Vector Institute for computing resources.
+We thank Nathan Samuel de Lara, James Ross, Hossein Goli, Qi Chen, Brandon Huang, Giseung Park, and others for helpful comments on drafts and conversations.
+Appendix Table of Contents
+- I Introduction
+- II Preliminaries and Related Work
+III Methodology
+- III-A Do Existing Diffusion Planners Compose by Default?
+- III-B Local Receptive Fields and Shift Equivariance in Score-Based Diffusion Models
+- III-C Additional Factors for Composition in Diffusion Planners
+- III-D Environment Descriptions
+- III-E Measuring Trajectory Diversity and Novelty
+- III-F Algorithm Setup
+IV Experiments
+- IV-A How Does Architecture Affect Composition?
+- IV-B Locality vs Shift Equivariance: How Much Do They Matter?
+- IV-C Replanning During Inference vs Architectural Design
+- IV-D Compositionality via Data Scaling vs via Architecture
+- IV-E Are More Compositional Models More Steerable?
+- IV-F Remaining Factors for Diffusion Planner Composition
+- V Discussion
+VI Appendix
+- VI-A More Related Work
+- VI-B More Method Details
+- VI-C Experiment Details
+- VI-D More Environment Details
+- VI-E Further Architecture Details and Experiments
+VI-F More Experiments
+- VI-F1 Do Common Diffusion Model Parameters Impact Compositionality?
+- VI-F2 Does Using the Continuous or Discrete SDE Formulation Impact Compositionality?
+- VI-F3 Does High-Temperature Sampling Impact Compositionality
+VI-A More Related Work
+Stitching through RL: Historically, stitching through bootstrapping has been seen as one of the primary benefits of temporal difference (TD) learning [45] in reinforcement learning (RL).
+Planning with diffusion models offers some of the stitching benefits of TD learning without needing to handle RL-specific challenges such as the deadly triad [39, 24, 73] and slow temporal credit assignment [12].
+Architecture in Diffusion Planning: Another work that similarly analyzes architecture and inference strategies for diffusion planning is Lu et al.
+[54], which interestingly finds that the best diffusion backbone are transformers, which have no inductive biases towards locality and usually have positional attention added.
+However, their work is about general performance in offline RL and not about emphasizing compositionality, which as discussed may not be an important properly for strong performance on the D4RL benchmark due to a lack of required compositionality [27].
+Diffusion Planner Inference Choices: When analyzing replanning, we analyze the most common form where replanning happens after a fixed number of execution steps.
+However, more sophisticated approaches to replanning exist which may have different effects on composition.
+For instance, Zhou et al.
+[81] adaptively re-plans when the estimated likelihood of the current partially executed trajectory falls too low.
+Other Forms of Compositionality: In our analysis of data scaling, we specifically analyze stitching as a property of data, but other works analyze other forms of generalization including concept composition [63], consistency [78] and grammar [25].
+Stitching in Non-Diffusion Imitation Learning: Prior to diffusion planners, other generative models were common in imitation learning [35].
+Most modern approaches are either based around autoregressive decoder-only transformers [68], such as the Decision Transformer [12], ACT [26].
+Other works analyzes stitching in imitation learning include these works [8, 34].
+We were inspired by Ghugare et al.
+[27], which analyzes stitching through a compositional generalization perspective.
 They revealed that applying supervised-learning as a form of offline RL [23] fails to stitch, requiring training data augmentation which manually stitches together trajectories.
 However, their result does not directly translate to ours because the properties we analyze (local receptiveness and shift equivariance) both enforce a form of regularization that creates composition, even when such behaviour is not explicitly built into the data.
-Compositionality in Diffusion Planners: Many approaches exist to create compositionality in diffusion planners.
-One common thread is performing hierarchical planning, including using diffusion to stitch existing dataset trajectories [41], generating subsequences jointly conditioned on each-other [48], or using a hierarchical cascade of models to separately create high-level plans and low-level trajectories [10, 42].
-Our work differs from these by mainly targeting whole-sequence compositionality, where composition does not occur from explicit stitching but from implicit model behaviour, similar to image diffusion.
-111See Appendix A.6.2 for more discussion.
-Several other works have observed that Diffusion models often produce stitched trajectories when frequently replanning during inference [9] and limiting the history the next generation is conditioned on [62].
+Other Works Examining Diffusion Planner Composition: Several other works have observed that Diffusion models often produce stitched trajectories when frequently replanning during inference [9] and limiting the history the next generation is conditioned on [72].
 Our work attempts to formalize these qualitative observations with more robust experiments and explain it under a single conceptual framework (locality for short history and shift equivariance for the replanning process) that also extends to architectural or training choices.
-Our focus on shift equivariance and locality was inspired by prior works which observed that these properties explain much of the compositional behaviour of image diffusion models [36, 52].
-We extend their work in a few significant ways.
-One, we connect their image-diffusion observations to sequential decision-making by connecting these properties that enable composition to other methods commonly used to achieve stitching in IL.
-Two, we show that common diffusion architectures fail to exhibit these properties, preventing stitching from occurring, especially at small data sizes.
-Three, we analyze the effect each independently has on compositionality.
-Finally, we show that common guidance strategies enable strong base-model compositionality to be directed towards certain goals, which is particularly important when using generative models for offline goal-conditioned RL.
-3 Methodology
-3.1 Local Receptiveness and shift equivariance in Score-Based Diffusion Models
-To unify prior explanations for compositionality in diffusion planners, we rely on two properties previously described for diffusion models in vision [36, 52].
-The first is locality, which means the size of the region around an individual state in a trajectory that influences the next denoising step for that state is constrained to a small region.
-A local model only attends to a small patch around part of a sample when denoising, but through multiple denoising steps, information can propagate across a generated sample.
-Definition 1: Locally Receptive Field.
-Assume we are interested in generating trjectories , and have a generative trajectory model whose output estimates the denoising gradient .
-We say that has a field locally receptive to a state neighbourhood if for all trajectories and state positions within those trajectories , .
-The second is positional equivariance, which requires that the model not use the absolute position of a state in the sequence as a condition during the denoising process.
-Convolutional neural networks (CNNs) are typically assumed to have this property, as convolutions by design shift a shared weight kernel and apply it to local patches in a sequence.
-However, downsampling and pooling operators common in CNNs break this property [70, 6].
-Definition 2: Shift equivariance.
-A trajectory score model is shift equivariant if for all shifts 222By shift equivariant we mean to discrete shifts and not continuous translations.
-See [49].
-on trajectories which uniformly change the position of states within the trajectory, .
-333In theory, circular padding is required for a CNN to have full shift equivariance[36], but in practice this is uncommon.
-3.2 Composition in Diffusion Planners
-Our main insight is that these notions of locality and shift equivariance have analogies to other strategies used to increase composition in diffusion planners.
-Increasing Compositionality by Replanning: The first strategy we address is replanning, specifically replanning with a limited history.
-A trajectory planning model replans when it generates a trajectory of length , executes some action horizon , and then generates a new trajectory starting at the current state .
-Because a diffusion planner can be queried to generate plans with a variable length, is often set to some number large enough to ensure temporal consistency far enough into the future to do effective planning.
-Additionally, a history of past states to condition on with length must be decided.
-Formally, given a generative trajectory model that generates trajectories of length conditioned on a history of length , each state in the trajectory will only be directly attentive to a local field of size , i.e, .
-It is important to note that locality through replanning and locality in the reverse-diffusion process are not identical, as the reverse diffusion process being several steps long means influence can be spread along a long trajectory through many local denoising steps.
-We show this visually in Figure 3.
-However, the similar mechanisms highlight how increasing the influence of nearby states in a given trajectory is important for composition.
-Replanning also ensures a form of shift equivariance when replanning with a fixed amount of the past trajectory kept as conditioning for the future trajectory.
+VI-B More Method Details
+VI-B1 Information Propagation Across a Sequence in Diffusion
+To visually illustrate how information can be propagated across a sequence in diffusion, we provide an illustration in Figure 12.
+VI-B2 How does shift equivariance work for diffusion policies?
 Diffusion planners can condition their plans on past state observations in two ways: either by including it as an explicit condition to the diffusion network, usually combined with classifier-free guidance (CFG) [37], or by inpainting states in the new generated trajectory.
@@
 Paradoxically, this means models with strong replanning capability must already have some form of shift equivariance.
-Increasing Compositionality by Training Augmentation: The next strategy to increase compositionality we address is training augmentation, where the data is modified during training to give the model approximate learned equivariance to some transformation [3, 59].
-The most common approach to encourage stitching is manually stitching together trajectories in the training set randomly by finding similar states in different trajectories[26, 8, 28, 39].
-This augmentation attempts to breaks the corelation between states far away in a trajectory that prevents stitching.
-However, this strategy requires tuning the distance metric that determines when segments are close, and is potentially infeasible if the combinatorial set of all legal stitches is large enough to make augmenting every possible trajectory expensive.
-We instead focus on augmenting the data to provide shift equivariance.
-Specifically, we examine positional augmentation, where the position of specific states within a trajectory are randomized during training such that the model cannot overfit to a state’s position.
-We implement this by sampling a random state from a training trajectory, removing the states before that state, and then padding the end of the trajectory with the final state until it is the same length as the original trajectory, such that the trajectory is randomly "shifted" by a number of states.444This assumes that the environment can be kept at the same state by the agent, which is true of our environments.
-Alternatives include masking out the remainder of the trajectory or allowing for mixed-length trajectories during training.
-Formally, given a training trajectory with states , we sample a position from the uniform distribution .
-To create our new augmented trajectory , we shift the states in by positions to the left, such that for state indexes , .
-For the remainder of the states , we set pad the states with the last state in the original trajectory: .This modified trajectory is then used when training our planner.
-Increasing Compositionality via Architecture Design Choices: The next strategy we address are architecture modifications.
-While equivariance can be learned through data augmentation, another popular strategy is to simply build equivariance to certain transformations through architecture changes[14, 15], especially in limited data settings.
-Surprisingly, we found that the common 1D U-Net proposed in [34] has a globally receptive field and is not equivariant to shifts, which we show in Appendix A.4.1.
-To easily analyze the effect that locality and shift equivariance have on composition, we replace the U-Net with a network which we call Eq-Net.
-Eq-Net simply consists of stacked 1D convolutions without pooling or down-sampling layers, which guarantees equivariance to shifts.
-We also decrease the size of the kernels in Eq-Net to guarantee locality.
-We defer further discussion and explanation to Appendix A.4.2.
-To test the effects of having a network with local receptiveness but not shift equivariance, we add positional encoding to Eq-Net with sinusoidal embeddings [66], adding each time positions’ embedding channel-wise after the first convolutional layer.
-See Appendix A.4.5 for details.
-Increasing Compositionality via Scaling Training Data: The last strategy we address is scaling training data, which has been frequently observed to increase diffusion model generalization [68] and compositionality [53].
-Importantly, this increase in compositionality does not come from memorizing more of the legal space of generations, but from learning the underlying rules governing composition in a particular area.
-Specifically, [23] finds that as data scales, progressively higher-level grammar rules are learned, starting with low level (local) consistency rules.
-Our work is specific to sequential decision making, where data is often sparse.
-For instance, the toy vision dataset CIFAR-10 [37] has samples, and the more modern ImageNet [17] has .
-By contrast, most of the environments in the modern offline RL OGBench dataset [54] have between and episodes per environment.
-Similarly, most of the tasks in [13] use or less trajectories.
-Given that most sequential decision making datasets have orders of magnitude fewer samples, considering other approaches to generalization becomes more important than relying on data scaling.
-Explaining Composition in Other Diffusion Planners: Our goal in examining the factors that lead to composition is to help explain why prior works compose (or fail to) so that future diffusion planning practitioners know which choices are important if their domain requires compositionality.
-Due to space constraints we include a thorough discussion of other works in Appendix A.6.3.
-Our main finding is that the majority of existing works achieve locality and shift equivariance either through frequent replanning with short-history conditioning [62, 9, 13] or some form of hierarchical decomposition where trajectories are generated explicitly as sub-trajectories and then composed manually [10, 48, 42].
-Surprisingly, achieving compositionality through architecture choices alone is rare, with our work to our knowledge being the first to take this approach in the planning domain.
-4 Experiments
-Our hypothesis is that local receptiveness and positional equivariance are key ingredients in explaining compositionality in diffusion planning, and that different ways of incorporating each (training augmentations, replanning, architecture choices, etc.) should increase compositionality.
-However, it is not known which of the two properties is more important, whether both are needed, or which way of inducting the two properties increases compositionality the most effectively.
-To test our hypothesis that local receptiveness and shift equivariance are sufficient for composition in diffusion planners and answer our open questions, we train a variety of models on toy environments designed to test generated sequence diversity and multi-goal completion.555Our code can be found here: https://github.com/rvl-lab-utoronto/diffusion-stitching
-Environment Description: Our main environment is a navigation task where an agent must navigate from one point in a grid to another while avoiding obstacles.
-Figure 1 shows examples of our grid environment, including points navigating to and from different locations in the grid.
-Additionally, to test training data scaling, we use a separate navigation environment where an agent moves between ten points in euclidean space in different sequences of points.
-More details are deferred to Appendix A.3.
-We chose toy navigation environments over complex benchmarks for a few of reasons.
-First, popular datasets which claim to test stitching either have testing start-goal pairs contained in the training set [24]666Shown by [26].
-or a very limited number of start and goal states [54].
-Additionally, existing datasets do not report what subsequences are present in the training data, making it difficult to predict stitching behaviour, whereas we can control this in our environment.
-Finally, the grid layout of our environment makes quantifying diversity easy, as each stitched trajectory will be topologically distinct [57], allowing us to perform careful analysis of a given model’s compositional ability.
-Data Collection for Diffusion Model Training: We look at the diversity of trajectories generated by diffusion planners in two separate planning settings, and collect two seperate datasets in our grid environment.
-The first is the fixed-endpoint setting, where the goal of the task is to reach a fixed end point from a fixed start point.
-Every training trajectory is optimal for the task, but the demonstration data includes multi-modal behaviour, as individual trajectories reach the end point via a different route.
-All trajectories have the same starting and ending state but take different paths through the grid, and the diffusion planner is tasked with generating new paths from the start to the goal composed of sub-trajectories seen in training in new configurations.
-The second is the end-point conditional setting, where a wider range of trajectories provide sub-trajectories that need to be composed to meet novel start-goal pairs.
-In our environment, the goal is defined as reaching a goal state from a start state, where each intersection point along the edge of the grid is a legitimate start or goal state.
-To be precise, we attempt to learn a model that can sample from the distribution .
-We collect data that includes diverse grid navigation to give the model ample sub-trajectories to compose, while being careful to keep some start-goal pairs unseen.
-We defer details to Appendix A.3.
-Algorithm Setup: For our experiments we use a 1D U-Net Diffusion Planner as the base model.
-We use CleanDiffuser’s [18] implementation and hyperparameters.
-One change we make is only predicting future states and using an inverse dynamics model to recover actions [1].
-777See Table 2 for details on hyperparameters and inverse dynamics..
-5 Results
-Here we report our findings on the unconditional and conditional environments, with the environments, data, and algorithms described in Section 4.
-In our results, we count diversity by examining topologically distinct trajectories.
-Practically, each trajectory can split the obstacles into two groups: those on its left side, and on its right; two trajectories with different obstacle splits are topologically unique.
-5.1 Unconditional Environment Results
-To quantify each model’s ability to compose in the unconditional environment, we took our final trained models and sampled them at a range of sample counts.
-From these samples, we counted what percentage of the possible (but unseen in training) trajectory space each model’s generations covered.
-By giving us a sense of how a model’s coverage scales with the number of generations, this gives a sense of how capable each model’s composition abilities are without solely relying on a model’s behaviour at very high sampling counts.888In our unconditional results, we filter to only analyze successful trajectories (ones that successfully reached the end state without hitting an obstacle) but all methods achieved very high success rates.
-Full completion rates are reported in Appendix A.2.1.
-Figure 4 shows our results for the unconditional navigation environment.
-Figure 4(a) shows a comparison of the different whole-sequence composition strategies.
-The base U-Net fails to ever compose on our small dataset, only reproducing examples seen in training.
-Our model with approximate positional equivariance (granted through training augmentation) performs better, but fails to extrapolate, not covering a significant proportion of the unseen trajectory space even at very high sampling rates.
-Our equivariant network with added positional encoding (making the model local, but not positionally equivariant) performs much better, indicating locality is a more important factor in encouraging composition than positional equivariance.
-Lastly, our model that is both locally receptive and shift equivariant (Eq-Net) performs the best, exhibiting very strong whole-sequence composition from architecture choices alone, without utilizing data scaling, augmentation, or replanning.
-Figure 4(b) analyzes replanning, where we progressively increase the number of times the model replans during a trajectory, and analyze the resulting diversity of generations over independent generations.
-For our replanning model we used our U-Net trained with positional data augmentation, as we found it necessary to enable the model to generate trajectories from diverse start positions.
-Assume we wish to replan times in a task long.
-When replanning, we keep only the current state ( using our prior notation), generate a new trajectory starting at that point, execute steps, and then repeat until the goal is reached.
-Figure 4(b) shows that as expected, as , the diversity of generations increases.
-To compare this to our architectural strategy, we compare it against the best model from Figure 4.
-To directly compare the computational cost associated with replanning, we compare each agent that replans times against Eq-Net that gets times as many parallel samples to generate from, such that each point queries the diffusion model an equal amount of times.
-For instance, at , the replanning model over independent samples is queried times, as each trajectory requires queries, so to have a fair comparison we query Eq-Net for independent samples.
-Predictably, our results show that more frequent replanning indeed leads to more trajectory diversity.
-Interestingly, they also show that the extra budget of diffusion model queries required during replanning might be better spent generating more samples from a high-diversity whole-trajectory model, as Eq-Net performs similarly or slightly better on all samples.
-Lastly, to validate that our results are not influenced by regularization induced by under-training of our diffusion planner, we analyze how each model’s compositionality evolves throughout training.
-We find that while early stopping helps compositionality (perhaps acting as a form of regularization that prevents overfitting to the training samples as has been observed in other work [5]), models with stronger inductive biases preserve their compositional abilities as they trained for longer.
-See Appendix A.5.1 for details.
-5.2 Conditional Environment Results
-In the conditional environment described in Section 4, the goal of the diffusion planner is to take unstructured trajectories in our environment and be able to consistently produce trajectories connecting an arbitrary start and goal location.
-To ensure that stitching is required at inference time, we examine every pair of edge intersections and filter out any pair where the starting pair and goal pair appear together sequentially, resulting in test-time pairs.
-We run our whole-sequence planners on each pair times, calculate an average completion rate and number of unique successful plans (where uniqueness is defined as topological distinctness), and take the average over sample runs.
-To guide the model to a specific goal, we use inpainting like in prior diffusion works [47, 34, 61].
-We found that repeating the inpainted goal state significantly improved goal-reaching performance, so we repeat the goal state times at the end of the trajectory before diffusing the intermediate states.
-Our results are shown in Figure 5.
-The model that is both local and positionally equivariant far outperforms all other models, achieving an average completion rate, roughly higher than the model with only locality but not positional equivariance.
-Only having positional equivariance did not substantially help the model in this scenario, with the U-Net with data augmentation only achieving a non-zero result on one of the testing pairs and only a small amount of the time.
-Lastly, like in the unconditional case, the default U-Net fails to perform stitching entirely if it is not locally receptive or shift equivariant.
-To visualize the difference in composition for each model, we plotted the unique successful runs from random generations from each algorithm on one of the hold-out start-goal pairs.
-Figure 6 shows the relative diversity of each model: the base U-Net fails to ever generate a successful trajectory, the U-Net made approximately positionally equivariant with data augmentation can produce a single candidate successful trajectory, the Eq-Net with positional equivariance removed produces a handful of plans, and the full Eq-Net produces a large diversity of successful trajectories.
-5.3 Data Scaling Results
-Lastly, we investigate the role of data in inducing composition.
-We use a different environment in these experiments, consisting of a navigation task where a point must sequentially navigate between five different points of a decagon.
-A compositional planner should be able to stitch sub-trajectories from individual points to other points.
-We train models with data ranging from from to data points, where points is close to the number of possible combinations.
-Figure 7 shows that even models without strong inductive biases can learn to compose at large enough data sizes.
-Past a certain data quantity (roughly samples in our environment), the U-Net learns to compose, producing new stitched trajectories at test time.
-We provide evidence that this data scaling induces a smaller receptive field of the diffusion model in Appendix A.5.2.
-This confirms that the ability to generalize composition of sub-sequences is another property that diffusion models learn from data at scale, adding onto concept composition [53], consistency [68] and grammar [23].
-6 Discussion
-6.1 What Are the Keys to Composition?
-While our work attempts explain why existing works in diffusion planning achieve composition, a reasonable question to ask is what engineering insights our findings have.
-While our results indicate that achieving whole-sequence compositionality through architecture choices may be under-explored, we emphasize that different choices are valid in different contexts.
-For instance, in environments where reactive control is very effective, generating long-horizon sequences may not be important, so frequent replanning with a short horizon may work better.
-In environments where data is abundant, using a higher-capacity model might be appropriate.
-We provide additional discussion of our findings in Appendix A.6.
-6.2 Limitations and Future Work
-Our work has several limitations.
-First, we only consider navigation environments with little complexity in the dynamics or observation space.
-While we motivate this in Section 5, evaluating our findings in a more realistic robotics environment would be useful future work.
-Second, our guidance strategy was limited to inpainting, which is effective in environments with clear endpoint goals, but less useful when goals are abstract (such as with language-conditioned agents) or multi-modal.
-Incorporating a way to leverage other guidance strategies like classifier-based [30] or free [31] guidance into our proposed high-compositionality diffusion planners is potentially fruitful future work.
-This would also let us fit our composition framework into the diffusion policy [13] setting, where composition of atomic skills may still be useful, but where a goal state cannot be inpainted because the diffusion model generates future actions and not future states.
-Appendix A Appendix
-Contents
-A.1 More Related Work
-Prior to diffusion planners, other generative models were common in imitation learning [29].
-Most modern approaches are either based around autoregressive decoder-only transformers [58], such as the Decision Transformer [12], ACT [25], or diffusion models [29].
-A more comprehensive survey on Diffusion Planners was prepared by [63].
-We draw from a rich body of work in diffusion model generalization, which typically examine image-based models.
-Approaches include examining the effect of the diffusion process noise [65], architectural biases [36, 52], and representation learning [35] among others.
-Another work that similarly analyzes architecture and inference strategies for diffusion planning is [46], which interestingly finds that the best diffusion backbone are transformers, which have no inductive biases towards locality and usually have positional attention added.
-However, their work is about general performance in offline RL and not about emphasizing compositionality, which as discussed may not be an important properly for strong performance on the D4RL benchmark due to a lack of required compositionality [26].
-Analyzing how transformers differ from CNNs with respect to compositionality remains unexplored as future work.
-When analyzing replanning, we apply a common form of it where replanning simply happens after some fixed number of execution steps.
-However, more sophisticated approaches to replanning exist which may have different effects on composition.
-For instance, [71] adaptively re-plans when the estimated likelihood of the current partially executed trajectory falls too low.
-A.2 Additional Experiment Details
-A.2.1 Unconditional Completion Rates
-| U-Net | U-Net+Aug | Eq-Net+PosEnc | Eq-Net | Replan (16) | |
-| Trials | |||||
-| Completion |
-Here we report completion rates in the unconditional environment.
-Table 1 shows the completion rates for each algorithm.
-As shown, every algorithm completes it close to of the time, with only the full-sequence models producing illegitimate trajectories a small portion (roughly ) of the time.
-A.2.2 Data Scaling Experiment Details
-For our data scaling experiments, we used a different environment, consisting of point navigation between points arranged in a decagon in euclidean space.
-Figure 8 shows a visualization.
-Each trajectory was tasked with travelling from point-to-point for a total of 5 visited points per trajectory.
-Data was collected with the same Gaussian noise as in the grid environment, and the trajectory length was also states long.
-We trained each of our models in this context for gradient steps, with hyperparameters otherwise the same as in Table 2.
-We ran each model times to approximate the distribution of the model’s output.
-The metric we report in 7 is the percentage of generations that are not seen in the training set, where membership in the training set is determined by finding which key-points a trajectory visits in which order, and comparing them to that particular model’s dataset.
-Note that this metric is somewhat biased towards the models that can compose but are trained on less data, as the space of compositionally legitimate but unseen in training trajectories is larger than for the model that already has most of the support of legitimate trajectories in its training set.
-However, our results show that the threshold is fairly binary, as the very low data models will not or will only very rarely produce a novel composition, but past a threshold the models will consistently produce novel trajectories a large percentage of the time.
-A.2.3 Algorithm Hyperparameters
-Here we report the common hyperparameter we ran our experiments with, unless otherwise specified. Table 2 shows the full values.
-| Hyperparameter | Value |
-| EMA Rate | 0.9999 |
-| Diffusion Steps | 100 |
-| Optimizer | AdamW [45] |
-| Base LR | 2e-4 |
-| Weight Decay | 1e-5 |
-| LR Schedule | Cosine |
-| Predict Noise or Clean | Clean |
-| Sampling Temperature | 0.5 |
-| Batch Size | 128 |
-| Gradient Steps | 500,000 |
-| Training Horizon Length | 512 |
-| Convolution Padding | Repeat |
-| Solver | DDPM[30] |
-| SDE Form | Discrete[30] |
-Instead of jointly generating states and actions as in [34], we use a separate inverse dynamics model (as in [1]).
+In practice however, because most diffusion policies are trained on a generation length much shorter than the horizon length (i.e, ) most of this equivariance comes from randomly sampling intermediate chunks of the trajectory during training.
+VI-C Experiment Details
+VI-C1 Baseline Descriptions
+For our baselines, we used the same core Diffusion Planning algorithm [40] with U-Net [40] and DiT [19] backbones, Diffusion Policy [13], and CompDiffuser [57].
+We used the implementations of Diffuser with U-Net and DiT and Diffusion Policy from CleanDiffuser [18] with minimal changes, and used our own implementation of CompDiffuser.
+Changes to each are elaborated on below:
+For Diffusion Policy, we condition on a stack of past proprioceptive observations but only on the current image. We did not perform any image augmentation to the images.
+For CompDiffuser, when conditioning on the starting or ending state, we repeated the state to match the overlap size between trajectories.
+We slightly modify their setup in training by first noising the prediction output, and then using the same randomly sampled diffusion step parameter to seperately noise the condition.
+We use exponential trajectory blending with on only the overlapping trajectory segments, like in the original paper.
+VI-C2 Algorithm Hyperparameters
+Here we report the common hyperparameters we ran our experiments with.
+We tuned each algorithm differently for the navigation environment (Maze) manipulation environments (the other five environments described in III-D) as we found the Maze’s more precise demands required different parameters to work well.
+For all environments, we down-sample the trajectory given by the diffusion model to a shorter horizon trajectory, sometimes called skip-planning [54].
+During inference we then linearly interpolate these down-sampled trajectories into a full-length one before they are given to the trajectory-following controller.
+We did this because we found that DiT on sequences longer than struggled with producing coherent trajectories and required unacceptable compute to train.
+For the navigation environments we use the same horizon across all algorithms; for the manipulation environments where fine-grained detail around picks was important we use less skip planning for the CNN-based models.
+We generally found the CNN models capable of handling planning horizons up to without issue.
+We performed min-max normalization to the observation and generation space for all algorithms, except the video inputs to visual Diffusion Policy, which we performed min-max normalization on.
+For Diffusion Planner and CompDiffuser, instead of jointly generating states and actions as in Janner et al.
+[40], we use a separate inverse dynamics model (as in Ajay et al.
+[1]).
 This removes the need to tune action versus state weighing in the diffusion objective, and generally performs better than joint prediction [54].
 We execute this inverse dynamics model with feedback, executing the action at each time step by giving the model the current state in the plan and the current state from the environment , such that .
-A.2.4 Experiment Computing Resources
-All experiments for this work were done on a laptop with a 13th Gen Intel Core i9-13950HX and NVIDIA GeForce RTX 4080 Laptop GPU.
-The final models used to produce the results shown in the paper were trained over roughly 2 weeks, and model inference to generate the final results took roughly 1 week.
-A.3 More Environment Details
-A.3.1 Grid Environment Details
-Our main environment is a small grid where a point agent navigates from one part of the grid to another while avoiding a series obstacles.
-The observation is the x-y coordinates of the agent, and the action is a x-y step which is normalized to a small step size.
-In the middle of the maze are several obstacles, which will kill the agent if it stays inside of the block for several steps in a row.
-Figure 1 shows examples of our environment, including points navigating to and from different locations in the grid.
-Navigating from one corner of the grid to the other requires steps out of a total allowed episode length of , challenging the diffusion planner’s long-horizon planning.
-The obstacles also ensure the trajectory must remain compositionally legitimate over the whole horizon, as accumulating error can cause the agent to wander off of the grid to its death.
-A.3.2 Grid Environment Possible Trajectories
+Full details for our Diffusion Planners are given in Table II, Diffusion Policy in Table III, and CompDiffuser in Table IV.
+[…]
+| Hyperparameter | Navigation | Manipulation |
+| Optimizer | AdamW [53] | AdamW [53] |
+| Base LR | 2e-4 | 2e-4 |
+| Weight Decay | 1e-5 | 1e-5 |
+| LR Schedule | Cosine | Cosine |
+| Diffusion Prediction | Noise | Clean |
+| Sampling Temperature | 0.5 | 0.5 |
+| Batch Size | 256 | 64 |
+| Convolution Padding | Repeat | Repeat |
+| Solver | DDPM | DDPM |
+| SDE Form | Continuous | Continuous |
+| Diffusion Steps Train | 100 | 100 |
+| Diffusion Steps Eval | 500 | 100 |
+[…]
+| Hyperparameter | Navigation | Manipulation |
+| Embedding Dimension | 64 | 64 |
+| Layer Sizes | (1,2,2,2) | (1,2,2,2) |
+| Planning Horizon | 128 | 512 |
+| Model Size | 120 MB | 120 MB |
+[…]
+| Hyperparameter | Navigation | Manipulation |
+| Embedding Dimension | 64 | 64 |
+| Model Dimension | 256 | 256 |
+| N Heads | 64 | 64 |
+| Depth | 8 | 8 |
+| Planning Horizon | 128 | 128 |
+| Model Size | 74.2 MB | 74.2 MB |
+[…]
+| Hyperparameter | Navigation | Manipulation |
+| Kernel Expansion Rate | 10 | 5 |
+| Channel Dimension | 256 | 128 |
+| Depth | 15 | 25 |
+| Planning Horizon | 128 | 512 |
+| Model Size | 61.1 MB | 48.9 MB |
+[…]
+| Hyperparameter | Maze | Well-Plate | Block-Stack |
+| To | 1 | 2 | 6 |
+| Tp | 16 | 16 | 16 |
+| Ta | 15 | 16 | 16 |
+| Img Res | N/A | (160,160) | (160,160) |
+| Video Encoder | N/A | ResNet18 | ResNet18 |
+| Gradient Steps | 1000000 | 1000000 | 1000000 |
+| Pos/Vel Control | Positional | Positional | Positional |
+| Architecture | Chi U-Net | Chi U-Net | Chi U-Net |
+| Learning Rate | 2e-4 | 1e-4 | 1e-4 |
+| Diffusion Prediction | Noise | Noise | Noise |
+| Diffusion Steps Train | 100 | 100 | 100 |
+| Diffusion Steps Eval | 100 | 100 | 100 |
+| Batch Size | 256 | 256 | 256 |
+[…]
+| Hyperparameter | Navigation | Manipulation |
+| High-Level Traj. Horizion | 128 | 512 |
+| Single Model Horizon | 32 | 128 |
+| # of Trajectories | 5 | 5 |
+| Gradient Steps | 1250000 | 1250000 |
+| Diffusion Prediction | Clean | Clean |
+| Architecture | U-Net | U-Net |
+| Batch Size | 250 | 250 |
+| Diffusion Steps Train | 100 | 100 |
+| Diffusion Steps Eval | 100 | 100 |
+| Inference Mode | Parallel | Parallel |
+VI-C3 Additional Architecture Experiment Details
+Here we show the completion rate of each architecture in our Maze environment.
+We found that completion dips slightly for Eq-Net near the boundry where it achieves compositionality, and a smaller dip for DiT where it achieves compositionality.
+See Figure 13.
+We consistently found that models experienced a dip in completion around the threshold they began to generalize (for Eq-Net around training samples, for DiT around training samples and for U-Net around training samples.
+Similar behaviour was seen in Pham et al.
+[67] where diffusion models undergo a three-phase transition as training data scales, starting with memorization, then undergoing a spurious generalization phase where sample quality drops, and finally achieving true generalization where quality is restored along with the achievement of creativity.
+VI-C4 Experiment Computing Resources
+Experiments used roughly 168 GPU-days of compute cluster use on GPUs equivalent to NVIDIA L40s.
+VI-D More Environment Details
+Here we describe our environments in more detail.
+Lights: We adapt this environment from the Lights Out/Puzzle environment from OgBench [64] with modifications.
+Specifically, in our environment pressing a button only changes the state of the button itself and not the surrounding buttons, and all trajectories start from the same state where no lights are pressed.
+We did both of these to make it easier to isolate when a generated trajectory has combined subskills from training trajectories.
+VI-D1 Dataset Details
+Table V describes details of the environment parameters and dataset description.
+[…]
+| Environment | Type | Embodiment | # of Trajectories | Trajectory Len | # of Obs | # of Acts | # of Possible Traj | # of Eval Traj |
+| Maze | Numpy | Point | 316 | 1000 | 2 | 2 | 705,421 | 10000 |
+| Didactic | Real | Franka | 3 | 12000 | 4 | 4 | 1 | 100 |
+| Cube-Stack | RoboSuite | Franka | 100 | 1000 | 4 | 4 | 720 | 500 |
+| Lights | OGBench | UR5e | 1000 | 500 | 4 | 4 | 65,536 | 2500 |
+| Well-Plate Sim | RoboSuite | Franka | 50 | 1000 | 4 | 4 | 7056 | 500 |
+| Well-Plate Real | Real | Franka | 50 | 51000 | 4 | 4 | 7056 | 100 |
+VI-D2 Data Collection
+Here we elaborate on our data collection process for each environment.
+Maze: In Maze, we collect data by starting with the point agent in the top left corner.
+At each intersection, it randomly selects whether to travel down or to the right at that intersection, and repeats for future intersections.
+If the point agent is along the bottom or right-hand face, it will only go to the right or down respectively.
+Well-Plate: In both the simulated and real Well-Plate environments, we enumerated all possibilities of start-plate cells to goal-plate cells (81, as there are 9 cells in each plate).
+Out of these 81 start-goal pairs, we randomly take 76 of them, and then randomly sample from the remaining pairs in groups of 3, excluding cases where there are collisions in either the start or goal plate.
+We did this until we had 50 trajectory descriptions consisting of three pick and places.
+To generate trajectories for these descriptions we generate a set of keypoints for the arm to travel between, and them follow them with a positional impedance controller.
+We fixed the RNG when generating the trajectory descriptions between the simulated and real well-plate environments, but the actual trajectories vary because of differences between the well plate placements between environments.
+Block-Stack: In Block-Stack, to generate a trajectory description we index the coloured blocks from 0 to 5, generate all ordered permutations of those indexes, and then randomly select 100 from the list.
+To execute description we use a similar system to Well-Plate where we generate keypoints guiding the arm to pick up each block and place it at the stack location, advancing the height the place occurs at in accordance with the growing height of the stack.
+Lights: We use the same trajectories in the puzzle-4x6-play-v0 dataset from Park et al. [64]. This environment has possible trajectories, as each button can either be turned on or off.
+Didactic: For Didactic, we manually wrote the set of keypoints for the three training trajectories.
+VI-D3 Environment Diversity Primitive Descriptions
+Here we describe the environment observation primitives we use to describe diversity, as defined in Section III-E.
+We also describe how many possible unique trajectories exist under our primitive formulation.
+We do not define primitives for Didactic, because the task is designed to only have possible trajectory that completes the task.
+Maze: In maze, the diversity primitives are the set of visited intersections in the maze.
+Intersections are the crossroads between the corners of four obstacles (or two obstacles and the wall).
+We calculate these visitations by taking the executed trajectory and seeing if the difference between each state in the trajectory and each intersection is smaller than some threshold.
+This means that each trajectory is defined by the set of intersections it passes through, and because each training trajectory moves from the top-left to bottom-right, each trajectory visits the same number of intersections.
+See Section VI-D5 for the number of possible trajectories.
+Well-Plate: In well-plate, the diversity primitives are the set of occupied cells in the goal plate, and the empty cells in the starting plate.
+Each training trajectory moves three blocks from the starting plate to the goal place, so each training trajectory is defined by a set of primitives.
+This makes the diversity measure blind to the order the blocks are moved in, but the end result is still a large space of potential trajectories.
+The total number of possible complete trajectories is equal to , as you choose 3 blocks from the starting plate to move to 3 cells in the goal plate.
+Block-Stack: In block-stack, the diversity primitives are tuples of a cube’s colour and it’s position along the stack.
+For instance, the set (Red,1),(Blue,2),(Yellow,3) is a possible trajectory description.
+The number of possible complete trajectories is , as you pick block from , then from the remaining , et cetera.
+Lights: In Lights, the diversity primitives are binary variables associated with each light being turned on (via being pushed).
+The number of possible trajectories is , as each light can either be turned on or off.
+VI-D4 Goal-Conditioned Environment Details
+For the Goal-conditioned Maze experiment, we generated a new dataset in the Maze consisting of random walks from points along the edge of the maze to other edge points with trajectories.
+We then found the set of all edge point start-goal pairs unseen in the same training trajectory to use as evaluation pairs.
+This ensure that to complete the unseen goals, the diffusion model must stitch sub-trajectories from the training set without simply replicating existing trajectories.
+VI-D5 Grid Environment Possible Trajectories
 Lemma 1: For a grid of size where is the number of blocks along the grid’s height and width, the number of possible topologically distinct trajectories that are optimal in length is upper-bounded (inclusive) by .
@@
 The size of this set is , because from positions we choose to be filled with either s or s.
-A.3.3 Data Collection
-In the unconditional setting, each trajectory starts in the top-left corner and navigates to the bottom-right corner, randomly switching from moving either downwards or right at each intersection.
-In total, our unconditional dataset includes unique trajectories out of a possible trajectories.999See Appendix A.3 for proof.
-This data quantity means our diffusion planner must generalize from a small number of samples to be able to generate samples in the (much larger) space of unseen but legitimate trajectories.
-In the conditional setting, we replicate the above data collection process, but for all four corners, to ensure a diverse set of subsequences.
-We additionally add data starting from each edge intersection point going towards other randomly selected edge intersection points.
-This ensures that the dataset contains a wide range of subsequences, but that the majority of start-goal states are not seen during training.
-This dataset contains trajectories.
-A.4 Further Architecture Details and Experiments
-A.4.1 U-Net is Non-Local and Not Shift Equivariant
-U-Net is Not Local: Here we support the claim made in the main text that the U-Net architecture proposed by [34] is not shift equivariant and displays heavy non-locality during the denoising process.
-To show this, we follow the procedure of prior works in analyzing diffusion model local field receptiveness by calculating the gradient of an intermediate sample point with respect to the input sequence.
-We took our final trained U-Net and Eq-Net models in the unconditional environment and generated samples.
-For each of the denoising steps, we calculated the gradient of the middle pixel in the sequence with respect to the whole input sequence (possible because of auto-differentiation), and averaged these gradients across the generations for each time step.
-Formally, for a diffusion model where is the next de-noised sample, is the previous sample, and is the denoising step, we calculate for each diffusion time-step from .
-Because we generate a 1-D sequence, we can plot these gradient maps as an image, where the x-axis is the denoising step and the y-axis represents the sequence gradient.
-Figure 9 shows these gradient maps.
-Visually, it is clear that the U-Net attends early in the denoising process to parts of the sequence far away from the centre pixel, even though the influence is much less strong than the pixels directly surrounding the centre.
-The size of the receptive field early in training spans far, containing non-zero information as far as the boundary states of the trajectory.
-On the other hand, Eq-Net has a hard boundary roughly pixels to each side of the centre pixel past which the model cannot attend to in a single denoising step.
-Like in prior works [36, 52], our diffusion models’ receptive fields progressively shrink during training, starting large and ending small.
-We also plot a rough estimate of the local receptive field size at each denoising step.
-We do this the same way as [52].
-We set a threshold value of the th percentile of the value of the absolute gradients in the image, find the window centred around the middle pixel that contains all values in that threshold for a given denoising step, and then plot the size of that window.
-Figure 10 shows the results of this, smoothed with a Savitzky-Golay filter.
-This plot shows that as expected, the local receptive field of the U-Net has a larger maximum value, approaching pixels wide at its max, while the Eq-Net is architecturally capped at .
-This is a somewhat surprising result, given that [34] specifically highlights the local receptiveness from convolutions as supporting composition.
-Given that our results in both the conditional and unconditional setting found that the U-Net architecture struggles with composition, we directly analyzed the network during the diffusion process to gain some insight into whether the model is sufficiently local.
-We suspect part of the reason for this is because while stride 1 convolutions are locally receptive if applied to the whole sequence, pooling operations and long-stride convolutions both down-sample the image substantially, such that the internal trajectory length at the bottle-neck of the U-Net may be small enough that even a small convolutional kernel is capable of attending to information far across the sequence.
-U-Net is Not Shift Equivariant
+VI-E Further Architecture Details and Experiments
+VI-E1 U-Net is Not Shift Equivariant
 Strictly speaking, empirical evidence that U-Net is not shift equivariant or that Eq-Net is should not be required, as it directly follows from known properties of their architecture: the U-Net has down-sampling strides which break shift equivariance [80], and Eq-Net has only 1-length strides with no pooling layers which preserves convolutions’ shift equivariance.
@@
 The U-Net architecture has substantial differences in the output when shifted, shown as a difference in the rightmost plot, while Eq-Net retains the same output but translated.
-A.4.2 Engineering Details for Eq-Net
-The main purpose of Eq-Net is not to propose a competitive architecture, but to show a proof of concept that locality and shift equivariance can be baked into architecture to enable stitching.
-Given this, we did not aggressively tune the architecture parameters of Eq-Net, instead finding values for network depth, kernel size, and channel size that worked well in our benchmark.
-However, we found it has some limitations and advantages that we will discuss here.
-Eq-Net is substantially more memory efficient than the base U-Net model, with the base U-Net having million parameters and Eq-Net only having million parameters.
-Despite this, Eq-Net was around half as efficient at both training and inference time.
-We suspect the former is because the channel sizes grew to be much larger in the intermediate representation layers of the U-Net ( at maximum) while Eq-Net has a fixed channel size of throughout.
-The latter is likely because Eq-Net is layers deep, which prevents efficient forwards passes because fewer operations can be parallelized.
+VI-E2 Engineering Details for Eq-Net
+For an architecture diagram of Eq-Net, see Figure 15.
+As noted in [42], in theory, circular padding is required for a CNN to have full shift equivariance, but in practice this is uncommon.
 Future work adapting these findings into a more practical architecture might investigate using adaptive polyphase sampling [6], anti-aliasing layers [80], or other modern tricks for incorporating shift equivariance into CNNs.
 Additionally, the depth of the network, way the layers of differing kernel sizes are layered, and channel size can likely be greatly optimized to increase computational efficiency and model capacity.
-A.4.3 Kernel Size Tuning
+VI-E3 Kernel Size Tuning
 The main parameter that requires tuning in Eq-Net is the kernel expansion rate (KER) which determines how fast the kernels grow deeper into the network.
@@
 Broadly, we found that tuning KER trades off between overall plan consistency (how likely the trajectories were to be locally consistent and stay outside the obstacles) and compositionality.
-To show this, we trained Eq-Net models in our unconditional setting with 3 KER values: 2, 5, and 12.
-Our results are shown in Table 3.
-We found that the intermediate value of 5 generated both a strong diversity of plans as well as reliably well-composed plans, while making the model too non-local made the model too local failed to consistently produce legitimate composed trajectories.
-| Large (KER = 2) | Medium (KER = 5) | Small (KER = 12) | |
-| Unique Trajectories | |||
-| Completion Rate |
-We also generated denoising gradient graphs of the small and large kernel Eq-Nets, like in our analysis of U-Net and our final Eq-Net architecture.
-Figure 13 shows the gradient graphs.
-As expected, Eq-Net with large kernels is attentive to information across the trajectory, and interestingly produces a gradient map extremely similar to the default U-Net, while Eq-Net with small kernels has an extremely limited receptive field.
-Our experiments with the default U-Net architecture and our Eq-Net architectures in total confirm our earlier hypothesis that while CNNs are meant to have strong inductive biases to encourage composition, the particular set of decisions made by [34] means at small data amounts they can still overfit heavily and attend to non-local informatin,
-A.4.4 Kernel Diversity
+To show this, we trained Eq-Net models in our unconditional setting with a variety of KER values.
+Our results are shown in Table VI.
+We found that the intermediate value of 10 generated both a strong diversity of plans as well as reliably well-composed plans, while making the model too non-local made the model too local and reduced completion rates.
+[…]
+| Task Completion % | 0.991 | 0.997 | 0.995 | 0.984 | 0.965 |
+| % of Generations Novel | 0.5056 | 0.752 | 0.802 | 0.869 | 0.9701 |
+Our experiments with the default U-Net architecture and our Eq-Net architectures in total confirm our earlier hypothesis that while CNNs are meant to have strong inductive biases to encourage composition, the particular set of decisions made by Janner et al.
+[40] means at small data amounts they can still overfit heavily and attend to non-local information.
+VI-E4 Kernel Diversity
 One thing that we found was important when designing Eq-Net was to include both very small (such as k=3) along with the larger kernel layers.
 When we used only medium sized kernels, it would lead to trajectories that were oddly "ungrounded" to the navigation plane, and would look correct in shape but would be shifted such that the resulting trajectory would nearly always hit obstacles.
-Figure 14 shows the visual results of this.
+Figure 16 shows the visual results of this in early experiments in a maze.
 We found that increasing the kernel size from small to large during training helped, somewhat mirroring the process in normal CNNs where fixed-size kernels stride over the whole image initially (capturing fine detail) and then smaller subsets after each downstride (capturing coarse detail).
 However, we did not explore whether the exact ordering of the kernel layers in Eq-Net matters or not.
-A.4.5 Positional Encoding
-When testing how our small receptive field model performs when it has access to positional information for each state, we use the same positional encodings as in [66], with slightly varied parameters.
+VI-E5 Positional Encoding
+When testing how our small receptive field model performs when it has access to positional information for each state, we use the same positional encodings as in Vaswani et al.
+[75], with slightly varied parameters.
 Figure 17 shows a visualization of how we add positional encoding to the Eq-Net architecture.
@@
 Specifically, the positional encodings are defined as
-We used a maximum length of instead of used by [66] because of the shorter length of our sequences.
+We used a maximum length of instead of used by Vaswani et al. [75] because of the shorter length of our sequences.
 It is worth noting that this strength of positional encoding (where the network is given enough information to precisely attend to the location of a state in the sequence) is likely quite a bit stronger than the positional attention from CNN aliasing.
 We included this as a representation of the worst-case scenario, to show the upper bound of models that are designed to attend to position, such as diffusion transformer architectures [66].
-A.5 More Experiments
-A.5.1 Overfitting Analysis
-Here we analyze the effect of overfitting.
+VI-F More Experiments
+VI-F1 Do Common Diffusion Model Parameters Impact Compositionality?
+In the following sections we report the findings of exploratory experiments we performed on a wide host of other diffusion model parameters we examined.
+Ultimately, we found that none of the following factors have significant effects on composition, with the exception of overfitting which has a moderate effect.
+All these experiments were done in our Maze environment, with reported statistics taken over evaluation samples.
+VI-F2 Does Using the Continuous or Discrete SDE Formulation Impact Compositionality?
+No.
+In diffusion models, the discrete and continuous formulations refer to different ways to parameterize the underlying stochastic differential equation governing the diffusion process [60, 7].
+Table VII shows that the continuous and discrete SDEs perform functionally identically.
+[…]
+| Completion % | 0.989 | 0.933 | 1 | 1 |
+| % of Generations Novel | 0.815 | 0.7894 | 0 | 0 |
+VI-F3 Does High-Temperature Sampling Impact Compositionality
+Surprisingly, not substantially.
+We expected the diffusion sampling temperature to have some effect, but instead found that models behaved similarly regardless of temperature, with some enhanced compositionality only occurring at temperatures high enough to greatly degrade sample quality.
+Table VI-F3 shows that high-temperature sampling does not make any model more creative.
+| Temperature | 0.1 | 0.5 | 1 | 2 | 4 |
+[…]
+| U-Net | 1 | 1 | 1 | 1 | 0 |
+| DiT | 0.991 | 0.991 | 0.995 | 0.952 | 0 |
+| Eq-Net | 0.989 | 0.987 | 0.989 | 0.993 | 0.26 |
+[…]
+| U-Net | 0 | 0 | 0 | 0 | NaN |
+| DiT | 0.0226 | 0.0221 | 0.0189 | 0.0297 | NaN |
+| Eq-Net | 0.827 | 0.801 | 0.806 | 0.815 | 0.796 |
+VI-F4 Does Predicting Noise vs Clean Impact Compositionality?
+No.
+In early experiments we found that predicting noise versus clean did not impact compositionality; however, it did impact sample quality (measured via completion rates).
+We found for the high-precision maze task that predicting noise worked better; when trained with the same parameters predicting a clean sample instead of noise, both U-Net and Eq-Net never successfully completed the maze.
+That our findings include environments where we train to predict noise (Maze) and where we train to predict clean (the manipulation ones) shows that the prediction target is likely irrelevant to compositionality.
+VI-F5 Does the Number of Diffusion Sampling Steps Impact Compositionality?
+Not past a certain, fairly low threshold.
+We found that while more sampling steps was very important to generate high-quality samples (measured by completion rate in our Maze environment, where slight perturbations from the maze causes the point agent to run into obstacles and die), we did not see a strong effect on compositionality, especially after roughly samples.
+Table VI-F5 shows the full results.
+| # of Diffusion Sampling Steps | 10 | 50 | 100 | 500 | 1000 |
+[…]
+| U-Net | 0.839 | 1 | 1 | 1 | 1 |
+| DiT | 0.409 | 0.914 | 0.973 | 0.994 | 0.995 |
+| Eq-Net | 0.105 | 0.807 | 0.927 | 0.989 | 0.992 |
+[…]
+| U-Net | 0 | 0 | 0 | 0 | 0 |
+| DiT | 0.009 | 0.00662 | 0.0258 | 0.009 | 0.032 |
+| Eq-Net | 0.688 | 0.796 | 0.813 | 0.804 | 0.81 |
+VI-F6 Does Overfitting Impact Compositionality?
+Yes.
 A common strategy in all machine learning to increase generalization is to perform early stopping, where training is ended before loss on the training set plateaus.
-Here, we run all of the models shown in the unconditional experiments at gradient increments of and run each model for samples, tracking both their completion rates and number of unique trajectories.
 Figure 18 shows that overfitting indeed reduces compositionality in the small-data setting, as models early in training achieve both high completion rates (indicating the generated plans are kinematically feasible) as well as more novel trajectories than later in training.
@@
 This hopefully indicates that the proper inductive biases from architecture choices can avoid the overfitting problem.
-A.5.2 Does Data Scaling Increase Locality?
-In the Results section, we found that scaling data increased the compositionality of our base U-Net model, even though it does not have a strong inductive bias towards locality.
-To see if the compositionality induced by data scaling is attributable to data increasing locality, we analyzed the models as done previously in Appendix A.4.1 by recording the gradient of the middle pixel for samples across each denoising step for each model.
-We then plot these directly, and also plot the estimated size of the local receptive field for each model.
-Following [52], to plot the receptive field we set a threshold value of the th percentile of the value of the absolute gradients in the image, found the window centred around the middle pixel that contains all values in that threshold for a given denoising step, and then plotted the size of that window.
-Figure 17 shows the gradient maps.
-Note that we name each data scale, with the amount of data corresponding to each name appearing in parentheses (SMM = Small Medium Medium, SM = Small Medium, SSM = Small Small Medium).
-Visually, it can be seen that the receptive field of the largest model (Full) is smaller than the field of the smallest model (Small) early in the denoising process.
-The small model attends to information clearly at the bounds of the image around step 20 of the denoising process, while the full data model does not have substantial gradient information past around 150 pixels away from the centre on either end.
-Figure 18 shows the plotted local receptive fields, smoothed with a Savitzky-Golay filter.
-The model trained on the highest data scale has a clearly smaller receptive field compared to the models trained on less data.
-This indicates that the reason why data scaling may help composition is because it trains the model to have a smaller receptive field, highlighting locality’s importance for composition and the usefulness of our theory.
-More work should be done to understand why data scaling specifically induces such a property.
-A.6 More Discussion
-A.6.1 Broader Impacts
+VI-G Expanded Main Text Experiment Results
+VI-G1 Data Scaling Gradient Maps
+Here we include the full gradient maps for the models described in Section IV-D. For spacing reasons, we include them at the very end of the appendix.
+VI-H More Discussion
+VI-H1 Broader Impacts
 Our work does not have any broader societal impacts beyond the general impacts that sequential-decision making research may have on automation or enabling more intelligent agents.
 While our work is limited to planning, some of our finding might generalize to diffusion models more broadly, which have implications for copyright/artist’s rights, deepfake generation, et cetera.
-A.6.2 How Should We Select a Method for Composition?
+VI-H2 How Should We Select a Method for Composition?
 While we focus on the whole-sequence generation case, it is not clear what the strongest method is to generate heavily temporally correlated, sequential data.
@@
 Hierarchical diffusion appears to be a strong way to incorporate both the benefits of long-horizon planning and short-horizon consistency.
-Besides the empirical benefits shown in the hierarchical diffusion planning papers discussed in 2, more evidence exists that hierarchical models can learn to minimize dynamics error in fewer training steps [19].
+Besides the empirical benefits shown in the hierarchical diffusion planning papers discussed in II, more evidence exists that hierarchical models can learn to minimize dynamics error in fewer training steps [20].
 One downside of these methods is that how the task is hierarchically decomposed either needs to be set manually, requiring more tuning, or must rely on a more complex algorithm [50] which increases implementation cost and other tuning.
@@
 However, recent work has found that taking a linear combination of the conditioned diffusion steps on a variety of past history lengths may preserve the benefits of both long-horizon planning and short-horizon composition, so it may be possible to combine the benefits of both [72].
-A.6.3 Explaining Composition in Other Diffusion Planning Works
+VI-H3 Explaining Composition in Other Diffusion Planning Works
 Here, we apply our framework to a couple of other works in the diffusion planning literature.
-Hierarchical Diffusion Planning: As discussed in Section 2, several works approach the composition problem by separately generating sub-sequences with a diffusion model, and then composing them.
+Hierarchical Diffusion Planning: As discussed in Section II, several works approach the composition problem by separately generating sub-sequences with a diffusion model, and then composing them.
 Exactly how these low-level subsequences are "orchestrated" into a single coherent plan varies, but each exhibits aspects of locality and shift equivariance.
+Chen et al.
 [10] propose Hierarchical Diffuser, which uses two separate diffusion models: a high level one which generates a "plan sketch", and a low level one which generates a sequence connecting the states in the plan sketch to "fill-in" the high-level plan.
-Functionally similar was the prior HDI paper from [42], which has a different form of generating sub-goals unrelated to diffusion.
+Functionally similar was the prior HDI paper from Li et al.
+[50], which has a different form of generating sub-goals unrelated to diffusion.
 These approaches enable a form of shift equivariance by data augmentation: because the low-level subsequences generated by the low-level diffusion model are randomly cut out of the longer trajectories in the training data, they have their relative position within longer sequences removed.
 This also enables locality by disentangling the joint prediction of states within a particular subsequence to any other states outside of it, so each subsequence functions as a small local window without connectivity to previous or future states, preventing confounding from preventing stitching.
+Chen et al.
 [10] also confirms some of our findings that Diffuser with a larger kernel size often produces more coherent long-term plans (4.3) but that this trades-off with worse ODD compositional generalization (Appendix E).
+Luo et al.
 [57] proposes a hierarchical method called CompDiffuser which, like the Hierarchical Diffuser paper, generates sub-sequences.
@@
 As discussed in the main text, short-memory and future prediction enforces locality by reducing the ability for long-horizon confounding effects to prevent composition.
-Results from [9] shows that this approach can enable compositionality, but only when keeping short or no memory and only generating short plans, essentially using the diffusion forcing model to approximate short-horizon Model Predictive Control (E.2).
+Results from Chen et al.
+[9] shows that this approach can enable compositionality, but only when keeping short or no memory and only generating short plans, essentially using the diffusion forcing model to approximate short-horizon Model Predictive Control (E.2).
 When extending the generation length or memory, they find that it instead reproduces existing trajectories.
-Follow up work from [62] extends Diffusion Forcing by replacing their RNN-based backbone with a transformer, and introducing the idea of combining the scores of the same model conditioned with several history lengths to try to achieve both long-sequence compositionality and short-term reactivity.
+Follow up work from Song et al.
+[72] extends Diffusion Forcing by replacing their RNN-based backbone with a transformer, and introducing the idea of combining the scores of the same model conditioned with several history lengths to try to achieve both long-sequence compositionality and short-term reactivity.
 They tested this approach on a task where a robot must take a fruit in a slot and move it to another slot.
@@
 Ultimately they found combining the score produced trajectories that both remembered important fast information and were capable of composing sub-trajectories for reactive control.
-This supports our argument highlighting the importance of locality for composition, but highlights that locality alone can run into issues in long-horizon tasks, as discussed in Section 6.1.
+This supports our argument highlighting the importance of locality for composition, but highlights that locality alone can run into issues in long-horizon tasks, as discussed in Section V-A.
 For details, see Sections 6.4: Task 3, C.8, and D.4.
@@
 Thus, in these models compositionality essentially comes from the same place that it does in TD learning: because no planning is done, the model composes through replanning.
-This replanning, like any explicit one-step or few-step policy, assumes the environment is Markovian and so does not include temporal information or preserve long history as conditioning, thus preserving shift equivariance and locality as we discussed in 3.2.
+This replanning, like any explicit one-step or few-step policy, assumes the environment is Markovian and so does not include temporal information or preserve long history as conditioning, thus preserving shift equivariance and locality as we discussed in III-C.
 There is some limited early evidence that when trained from scratch on very small datasets, Diffusion Policies may perform learning more similar to memorization with nearest-neighbour interpolation than "generalization" in any meaningful sense, and thus obtain strong apparent trajectory composition just through being very reactive [32].
 However, these results have only been found in very small data settings, so it is likely that they do not apply to larger VLAs or to multi-task Diffusion Policies trained on a heterogeneous mix of tasks.
+VI-I What explain Composition in Prior Monolithic Diffusion Planners?
+Our results shown in Section III-A and throughout the paper indicate that diffusion planners, when used to generate long-horizon plans, tend to memorize exact sequences.
+This seems to contradict the results of Janner et al.
+[40] and Ajay et al.
+[1] which both also use monolithic planners.
+We suspect the following reasons are at play:
+Most of their tasks are trained on the order of thousands of trajectories, which we find is usually sufficent to generalize.
+The D4RL Maze datasets have been found not to require stitching [27], potentially letting the diffusion planner "cheat" on the task by simply memorizing appropriate trajectories.
+For many of their experiments they use a shorter planning horizon of , allowing some stitching from replanning like with diffusion policies [13].